### v5

11institutetext: Romina Etezadi 22institutetext: Sallam Abualhaija 33institutetext: Chetan Arora 44institutetext: Lionel Briand 55institutetext: 55email: retez068@uottawa.ca, sallam.abualhaija@uni.lu, chetan.arora@monash.edu,
lbriand@uottawa.ca
1 School of Electrical Engineering & Computer Science, University of Ottawa, Canada
2 SnT Centre for Security, Reliability, and Trust, University of Luxembourg, Luxembourg
3 Faculty of Information Technology, Monash University, Melbourne, Australia
4 Lero Research Ireland centre for Software Research and University of Limerick, Ireland
Classification or Prompting: A Case Study on Legal Requirements Traceability
Romina Etezadi1    Sallam Abualhaija2    Chetan Arora3    Lionel Briand1,4
1 Introduction
Technological advancements are significantly transforming software development across diverse domains, such as healthcare Caruana:15 . Software applications and automated assistants have become integral to our daily lives Zhan:22 . This evolution, driven by recent breakthroughs in artificial intelligence (AI), has led to increasing complexity in software systems ahmad2023requirements ; Feldt:18 . As technology progresses, regulations are adapting in parallel to ensure that software systems are developed in line with ethical and legal standards. For example, the general data protection regulation (GDPR) GDPR is enforced since 2018 to address concerns about privacy and data protection. Despite being introduced by the European Union (EU), the GDPR has a global effect, impacting organizations (and software) outside the EU as long as they handle personal data of EU residents.
Requirements Engineering (RE) plays a pivotal role in this landscape. RE is concerned with specifying and maintaining software requirements that outline the properties and functions of a system-to-be Pohl:11 . Legal compliance of software systems against applicable provisions can be addressed at different stages of software development. One scenario is to explicitly identify legal requirements early during the requirements elicitation phase, answering the question: “What legal obligations need to be satisfied by the system for it to be compliant?”. The elicited legal requirements can then be integrated into the software development process, while maintaining trace links to the source legal provisions. As an alternative scenario, requirements engineers may need to verify the compliance of existing software systems against legal provisions in a post-deployment stage, as new regulations have become applicable. In this case, they must answer the question “Does the system satisfy the regulation?”. To do so, engineers must analyze the regulation, identify applicable legal provisions, and then trace software requirements to these statements. Both alternatives rely on requirements traceability analysis, an essential RE activity concerned with the identification and maintenance of trace links between requirements and other artifacts within the software development lifecycle Meyer:22 . legal requirements traceability (LRT) is a special case where requirements are traced to provisions in a regulation and is the focus of this paper.
To illustrate this concept, consider the following example. Imagine a fictional mobility app named WeMobilize, which helps users book and share cab rides. Originally a non-EU startup, WeMobilize is expanding to the EU and hence must comply with the GDPR. This example is particularly relevant as many businesses are globalizing and must adapt to data protection laws in different jurisdictions. Fig. 1 shows how WeMobilize’s requirements (labeled REQ1 – REQ5) can be traced to data protection policies in the GDPR GDPR . We identify trace links to provisions in GDPR for REQ1 and REQ3 – REQ5, visualized as dashed lines in black. REQ2 has no trace link to GDPR in our example since it does not involve processing users’ personal data.
REQ1 involves collecting user’s personal information and must therefore be traced to two provisions, namely REG_DIR (related to the direct collection of personal information) and REG_CON (related to the explicit soliciting of users’ consent). Currently, consent is not part of REQ1, which prevents identifying a trace link with REG_CON—a missing trace link is visualized with a red dashed line in the figure. Failing to identify this trace link entails a possible breach of GDPR. Therefore, deploying WeMobilize as-is, without accounting for provisions in GDPR, can lead to potential reputational and financial losses caused by violating the GDPR. LRT can help identify potential non-compliance issues at early stages but requires not only legal expertise but also substantial manual effort. Developing automated support is therefore beneficial to assist engineers and analysts in identifying applicable trace links.
Figure 1: Example on tracing WeMobilize app requirements to GDPR statements.
However, achieving this is challenging for LRT due to: 1) Language Discrepancy: There is a mismatch in vocabulary, style, structure, and abstraction level between legal regulations and software requirements. Legal text often relies on legal terminology and remains intentionally generic, while software requirements are often written in technical or domain-specific language. This discrepancy must be carefully considered when selecting or designing LRT solutions. For instance, REQ4 states, “I want to stop receiving notifications.” Although the term “consent” does not appear explicitly, a human can infer—using common sense and legal understanding—that the ability to stop receiving communications relates to the right to withdraw consent (REG_WCON). Methods that cannot bridge this conceptual gap may fail to establish such trace links. This highlights the need for models that can reason across domains. 2) Limited Training Data: Another major challenge is the scarcity of annotated data for the LRT task. Bridging language discrepancies often requires training modern NLP models on domain-specific examples to effectively capture cross-domain semantics. However, in industrial settings, obtaining large volumes of traced requirement-regulation pairs is difficult due to privacy, confidentiality, or limited documentation. Due to the scarcity of training data, it is crucial to prioritize selecting or designing models that perform robustly with limited examples.
Requirements traceability is a well-explored problem in the RE literature, e.g., wang2018requirements ; tufail2017systematic . However, the extensive research on requirements traceability is not directly applicable to LRT due to the challenges outlined above. Despite the serious consequences of non-compliance, LRT has received limited attention from the community. Cleland-Huang et al. cleland:2010 ; Gibiec:2010 proposed a classifier that predicts trace links by computing the likelihood of requirements being traced to provisions based on indicator terms found in both provisions and requirements. Guo et al. Guo:17 focused on bridging the terminology gap between provisions and software requirements. They examined three methods, including the one by Cleland-Huang et al. mentioned above, and two others based on web-mining and ontologies. The proposed methods aim to expand the terminology of the provisions by adding additional terms to better identify trace links.
Existing traceability approaches for the LRT task, mentioned above, have several limitations. First, they leverage a statistical model based on TF-IDF (Term Frequency-Inverse Document Frequency), which primarily relies on lexical overlap between documents, where similarity is computed from the presence and frequency of overlapping words. This approach is mostly effective when the source and target texts share similar vocabulary. Still, they struggle to capture deeper semantic relationships, which are required in LRT, as the language used in legal texts often differs significantly from that in technical requirements. Second, existing techniques have yet to fully leverage the recent advances in natural language processing (NLP), evolving from latent semantic methods such as Latent Semantic Indexing (LSI) deerwester1990indexing and Latent Dirichlet Allocation (LDA) blei2003latent , which uncover hidden semantic structures through concept and topic modeling, toward more contextually rich approaches. We posit that (large) language models (LMs or LLMs) and contextual sentence embedding architectures offer capabilities beyond lexical or shallow syntactic matching, enabling them to capture nuanced legal and technical semantics crucial to LRT. While some recent work has applied language models to traceability tasks lin2021traceability , these efforts have not been tailored to the distinctive demands of aligning with legal requirements—such as interpreting domain-specific terminology, resolving cross-references, or bridging differences in stylistic and structural conventions between legal and technical texts. Architectures like Sentence Transformers can generate semantically rich representations that are better aligned with these needs but remain underutilized in this context. Third, prompt engineering has not been given sufficient attention; existing work often relies on overly simple prompts ronanki2024requirements ; hey2025requirements ; ge2025cross or on querying multiple LLMs fuchss2025beyond , which may not always be feasible in real-world deployments due to computational and cost constraints. Fourth, the evaluation is based on a single benchmark that may not fully reflect the complexity of the legal domain in practice. To address these limitations, we propose two novel approaches in this paper, based on recent NLP technologies that utilize the Transformer architecture Vaswani:17 and LLMs. Similar to existing work, both approaches aim to predict trace links, and we assess their performance on a realistic scenario beyond the commonly used benchmark dataset.
Contributions. The paper makes the following contributions:
(1) We devise two automated approaches (Section 3) for predicting trace links between requirements and provisions based on LLMs. Our first approach, hereafter referred to as $\mathcal{K}$ashif, standing for automated trace linK identificAtion between legal proviSions and tecHnical requIrements using sentence transFormers. $\mathcal{K}$ashif leverages sentence transformers (STs), pre-trained language models optimized for understanding longer text sequences such as sentences, and predicts trace links based on semantic similarity. STs are generally more effective than word-level language models Reimers:19 for tasks like LRT because they capture the holistic semantic meaning of entire sentences rather than relying on individual word representations. While word-level models like BERT, in their default form, produce contextual embeddings for words and require post-processing (e.g., averaging) to represent sentences, they can also produce sentence-level representations through their pooled output (e.g., the [CLS] token) for classification tasks. STs, on the other hand, are a modification of the word-level models that employ siamese and triplet architectures to generate semantically meaningful sentence embeddings Reimers:19 , and are explicitly trained on sentence pairs using objectives like contrastive loss. This allows them to learn semantic similarity at the sentence level, making them better suited to identifying conceptually equivalent statements—an essential capability in LRT, where the legal and technical texts often have minimal lexical overlap. Our second approach utilizes Rice_LRT, a recent framework that enables effective prompting of LLMs. We employ Rice_LRT with the GPT-4o model offered by OpenAI [^1] https://openai.com/index/hello-gpt-4o/. Our solutions are described in Section 3.
(2) We empirically evaluate our first solution, $\mathcal{K}$ashif, on a benchmark dataset, referred to as HIPAA Guo:17 , comprising textual requirements traced to 10 different provisions. We further compare $\mathcal{K}$ashif against a baseline classifier from the literature cleland:2010 ; Guo:17 . Moreover, we compare $\mathcal{K}$ashif with five other baselines employing different technologies, as fully explained in Section 4.3. We re-use, re-implemented, and re-evaluated the baselines as part of this work. Our evaluation shows that $\mathcal{K}$ashif yields an average F2 score of $\approx$63%, leading to a substantial improvement of 34 percentage points (pp) over the best baseline. While $\mathcal{K}$ashif still performs significantly better than the baselines, such accuracy is rarely practically useful in real-life scenarios where the number of provisions easily exceeds 10 (as is the case in HIPAA). More details on this evaluation can be found in Section 4.6.
(3) To further confirm its performance, we test $\mathcal{K}$ashif on new unseen requirements documents covering diverse domains and requirements types. These requirements are traced to the GDPR, a more complex regulation with 26 provisions pertaining to personal data protection that must be adhered to in software requirements. On this dataset, the base ST model (without additional fine-tuning) yields an average recall of 15%. In comparison, a pre-trained sentence transformer, with no exposure to the requirements traceability task, yields a nearly zero recall, as elaborated in Section 4.7. Driven by this observation, we propose our second solution, the final contribution of this paper, as explained next.
(4) We devise a prompt strategy based on the Rice framework, capturing recent state-of-the-practice in LLMs for RE. For simplicity, we refer to our prompt strategy hereafter as Rice_LRT. Our evaluation (reported in Section 4.8) shows that using Rice_LRT with the GPT-4o LLM leads to an average accuracy of 84% in predicting trace links in the GDPR dataset, a complex and general regulation. Compared to $\mathcal{K}$ashif, Rice_LRT shows a remarkable gain of 69 pp in accuracy. Moreover, compared to simpler prompts, our engineered prompt outperformed them by 34 pp. Using Rice_LRT in practice can significantly reduce the time and effort required to manually identify trace links. With Rice_LRT, the analyst will vet only a small fraction of the provisions, equivalent to $\approx$9.5%, while identifying 84% of actual trace links. Further, GPT-4o also provides an informative rationale for each predicted trace link.
Structure. Section 2 provides background. Section 3 presents our proposed approaches. Section 4 reports on our empirical evaluation. Section 5 discusses threats to validity. Section 6 reviews the related work, and finally, Section 7 concludes the paper.
2 Background
Language Models (LMs). Language Modeling in NLP involves computationally determining the probability distribution of word sequences Jurafsky:20 . Given a sequence of words, an LM predicts the most likely next word, enabling it to generate text Alexandrescu:06 . For example, an LM would predict “Mat” as the most likely next word in the input sequence, “The cat sits on the [WORD]”. LMs are trained on large corpora of texts to estimate these probability distributions accurately. State-of-the-art LMs are based on transformer architecture, which leverages self-attention mechanisms to weigh the significance of different parts of an input text relative to a given position Vaswani:17 . The attention mechanism determines which words in a sentence are more important based on the context and gives them more “attention”. For instance, in the sentence “Mary, who used to live in Paris, loves wine.”, the attention is on Mary and wine. Building on transformer architectures, the Sentence Transformers framework (ST) Reimers:19 offers a set of pre-trained models designed to encode longer text sequences, such as sentences or paragraphs, into dense vector representations within a high-dimensional space. They produce contextual embeddings that capture the overall semantic essence of an entire input sequence.
More recently, generative LLMs have emerged as transformer-based language models that are scaled up significantly in model size and the amount of training data. Examples on LLMs include OpenAI’s GPT (Generative Pre-trained Transformer) Radford:18 and LLaMa touvron2023llama ; touvron2023llama2 . These models can perform new tasks based on textual instructions (prompts) NEURIPS2020_1457c0d6 .
Machine learning (ML). Supervised learning is one of the most prominent paradigms in ML. In this paradigm, the ML algorithm is provided with labeled training data where each data point consists of an input vector (features) and the corresponding output label (or value). The algorithm learns patterns within the input features to make predictions based on this training. When trained on a sufficiently large dataset, the algorithm refines its predictions to classify the provided labels more accurately. The example ML classification algorithms include random forest, decision tree, support vector machine, and feed-forward neural networks Goodfellow:16 .
3 Approaches
This section defines our notation and then presents our proposed approaches, $\mathcal{K}$ashif and Rice_LRT.
3.1 Notation
[…]
To predict trace links between requirements and provisions, $\mathcal{K}$ashif utilizes Sentence Transformers (ST) and cosine similarity Manning:08 .
3.2 $\mathcal{K}$ashif
Fig. 2 provides a comprehensive overview of the two phases comprising our approach. Phase A covers steps 1-3 and offers a developer’s perspective, focusing on building a traceability model for solving LRT. Step 1 prepares a training dataset of manually identified trace links. Step 2 selects a pre-trained model to customize for addressing LRT. Step 3 involves fine-tuning the LRT model. Phase B covers steps 4-6 and provides the perspective of an end user (e.g., a requirements analyst) assuming the availability of an LRT model. Step 4 preprocesses the input requirements document (RD). Step 5 applies the LRT model to compute the semantic similarities between each requirement in the RD and each provision. Step 6 predicts trace links. We explain these steps in detail next.
Figure 2: Overview of $\mathcal{K}$ashif.
Step 1: Training set preparation
Step 1 assumes the availability of a labeled dataset for LRT. We discuss the dataset used in our work in Section 4.2. In this step, we transform the training examples into a format suitable for fine-tuning the pre-trained ST models. Each training example is represented as a triple $\langle r_{i},c_{j},\ell\rangle$, where $\ell=1$ when $r_{i}$ and $c_{j}$ have a trace link (positive sample) and $\ell=0$ (negative sample) otherwise.
Step 2: Model Selection
Defining which pre-trained models to start with has become a challenging task due to the regular release of new models [^2] As of May 15, 2024, there are 124 ST pre-trained models available on HuggingFace.. Ideally, one should fine-tune all available models to select the best-performing one. However, since fine-tuning is resource-intensive, we narrow down the alternatives for experimentation in this step. Selecting the best ST model in step 2 is the subject of RQ1, elaborated in Section 4.5.
Step 3: Model fine-tuning
[…]
The resulting LRT model is then passed on to step 5.
Step 4: Preprocessing
In step 4, we preprocess the input requirements using a simple NLP pipeline composed of two modules, namely Tokenization and sentence splitting. The goal is to decompose the text into separate sentences. In our work, a requirement $r_{i}$ corresponds to a sentence generated by the NLP pipeline, which may or may not be grammatically correct. Using $\mathcal{K}$ashif to solve LRT for multi-sentence requirements is straightforward. A provision is traced to the requirement if it is traced to any sentence thereof. The intermediary output of this step is a set of $n$ requirements ($\mathcal{R}=\{r_{1},r_{2},\ldots,r_{n}\}$) from the input RD.
Step 5: Similarity Computation
Given a set of $m$ provisions $\mathcal{C}$, step 5 computes the semantic similarity scores between each $r_{i}\in\mathcal{R}$ and each provision $c_{j}\in\mathcal{C}$. In this work, we apply cosine similarity, which is a widely-used measure for text similarity Jurafsky:20 . The similarity score is a real value between 0 to 1. A score close to 0 indicates dissimilarity, while a score close to 1 indicates similarity. The output of this step is a matrix of dimension $n\times m$, containing the similarity scores between the $n$ requirements in the RD and the $m$ provisions in $\mathcal{C}$.
Step 6: Trace links Prediction
Step 6 predicts a trace link between $r_{i}$ and $c_{j}$ using the similarity matrix from step 5. A trace link is predicted when the similarity between $r_{i}$ and $c_{j}$ exceeds a certain threshold $\theta$. Below, we discuss alternative methods for setting $\theta$.
(a) Constant Threshold: To predict a trace link, we utilize a pre-defined constant threshold, $\theta=0.5$. Specifically, a trace link is predicted if the similarity score exceeds $0.5$. This threshold is considered a reasonable rule of thumb, as evidenced by its previous application in the literature Yao2014 ; Corley2005 . Moreover, this threshold was chosen based on the fine-tuning approach used in $\mathcal{K}$ashif. During training, positive pairs are labeled with a similarity score of 1, while negative pairs are labeled with 0. The model is thus optimized to produce similarity scores closer to 1 for true links and closer to 0 for unrelated pairs. As a result, a threshold of 0.5 naturally serves as a midpoint between these two extremes, making it a natural choice for distinguishing between positive and negative links.
(b) Dynamic Threshold: Another practical method to adjust $\theta$ involves curating a set of negative training examples, i.e., requirements that do not have trace links. These requirements can be sourced from publicly available datasets or from different projects. However, for more accurate results, it is ideal to use requirements from the same project under analysis. Inspired by similarity-based classification proposed in the literature Amaral:21 , we select $\theta$ using the following procedure. For each provision $c_{j}\in\mathcal{C}$, we identify a set of negative training examples ($TR_{j}^{-}$), i.e., requirements $\{r_{1}^{\prime},\ldots,r_{k}^{\prime}\}$ that do not have trace links to $c_{j}$. We then compute the similarity between $r_{i}$ and $TR_{j}^{-}$ and set $\theta$ to the average cosine similarity between $r_{i}$ and $TR_{j}^{-}$. If the similarity between $r_{i}$ and $c_{j}$ is higher than the similarity between $r_{i}$ and $TR_{j}^{-}$, then $r_{i}$ is semantically closer to $c_{j}$ and should be traced to it. Conversely, if the similarity between $r_{i}$ and $TR_{j}^{-}$ is higher, then it should not be traced to $c_{j}$ as it is semantically closer to the negative examples. This procedure sets a different $\theta$ value for each $r_{i}$ based on randomly selected negative examples.
(c) Maximum Delta Cutoff: In this method, we apply the following procedure. First, for each $r_{i}$, we sort the similarity values computed across the different provision $c_{j}\in\mathcal{C}$. Then, we compute delta values ($\Delta$) corresponding to the differences between each pair of consecutive similarity values and identify the largest $\Delta$ (i.e., the biggest gap in the computed similarities). To illustrate, consider the following example. Assume $r_{i}$ has similarity values of 0.98, 0.1, 0.3, and 0.7 with $c_{1}$, $c_{2}$, $c_{3}$, and $c_{4}$. We sort these values in descending order as follows: $c_{1}$: 0.98, $c_{4}$: 0.7, $c_{3}$: 0.3, $c_{2}$: 0.1. Next, we compute the $\Delta$ values: $\Delta(c_{1},c_{4})$=0.28, $\Delta(c_{4},c_{3})$=0.4, $\Delta(c_{3},c_{2})$=0.2. Based on these values, the largest $\Delta$ is 0.4 between $c_{3}$ and $c_{4}$. Finally, we set $\theta$ to the lower similarity value in the pair that yielded the largest $\Delta$. In the above example, we would set $\theta$ to 0.3 (the similarity value between $r_{i}$ and $c_{3}$). The largest $\Delta$ represents the most significant drop in similarity, indicating a potential boundary between relevant and irrelevant provision for $r_{i}$.
(d) Tuned: In this variation, rather than using a fixed threshold $\theta$, we optimize $\theta$ by performing a search over the training set, evaluating values in the range [0.01, 1) with a step size of 0.01. The optimal threshold with the highest F2 score identified on the training set is then applied to the test set.
The methods described above result in four variants of $\mathcal{K}$ashif, each determined by how $\theta$ is set. These variants are referred to as $\mathcal{K}$ashif constant, $\mathcal{K}$ashif dynamic, $\mathcal{K}$ashif Δ, and $\mathcal{K}$ashif tuned. We compare these variants in Section 4.
3.3 Rice_LRT
Our second proposed approach, Rice_LRT, is composed of two steps as illustrated in Fig. 3. The first step involves designing a prompt that is effective for addressing LRT. The second step then applies the prompt to instruct an LLM to predict trace links. We elaborate on these steps next.
Figure 3: Overview of Rice_LRT.
Step 1: Prompt Design
In this step, we designed the prompt following recent best practices reported in the RE literature vogelsang2024using ; vogelsang2024specifications . Fig. 4 presents our final prompt, obtained through iterative refinements. The prompt was used on each requirement in the input RD, with the requirement located at the end of the prompt. To design the prompt, we followed the Rice (Role, Instruction, Context, Constraints, Examples) framework, with some small adaptation to the LRT task, as we discuss below. The prompt is structured in the following five elements:
  • Context: This element introduces the LRT task. Since the role is implicitly indicated as a requirements analyst building the trace links, this element subsumes the Role element in the original Rice framework and simply provides the Context. We omitted the explicit mention of the role to obtain a more general applicability of the prompt. For LRT, multiple analysts with different backgrounds are likely involved, e.g., a legal analyst in addition to the requirements analyst. The context corresponds to the text shaded in cyan in Fig. 4.
  • Examples: This element provides a few examples selected from our ground truth. The examples should cover different trace links. Each example is composed of a requirement and the set of trace links alongside the rationale behind each trace link. We note that the LRT task is complex, as we demonstrate throughout the paper. For this reason, we opted for the few-shot prompting technique. This element matches Examples in Rice. Examples corresponds to the text shaded in pink in Fig. 4.
  • Instruction: This element provides explicit instructions on how to perform the LRT task. This element aims to guide the model through the right reasoning process to generate the desired output. The Instruction element corresponds to the text shaded in olive green in Fig. 4. Compared to the original Rice framework, this element contains both the Instruction element combined with the Constraint. The reason for this is that both elements are intertwined in our context. The prompt must therefore account for task-specific considerations, explained below.
    • $\bullet$
      The prompt should encourage the LLM to equally consider other provisions, since only a subset of the provisions are explicitly explained via the examples and rationales in the Examples element. Ideally, the prompt should present an example on each provision. However, this is infeasible since only relevant provisions should be traced to software requirements in a given project. For instance, if the legal basis for collecting personal data is the contract, then unlike explicit consent, only certain data subject rights are applicable according to GDPR and must be appropriately implemented in the software.
    • $\bullet$
      The prompt should account for indirect trace links. As stated above, the LRT is challenging primarily due to the terminology gap between requirements and provisions. We therefore encourage the LLM to use its reasoning capability to identify indirect links, generalizing beyond the provided examples in the prompt.
    • $\bullet$
      The prompt should favor recall by predicting at least one trace link for each requirement. As we discuss in Section 4, filtering out falsely introduced trace links, as long as they are not too numerous, requires less time and effort by the human analyst than identifying missing trace links.
  • Output Indicator: This element clearly describes the output format, corresponding to the text shaded in violet in Fig. 4.
We note that prompt templates such as Rice serve as illustrative examples or starting points rather than scientifically validated or universally optimal configurations vogelsang2024specifications ; huang2025prompt . Prompt templates are typically task-specific and heuristic in nature. Based on our preliminary analysis of Rice, providing the example before instructions and contextualizing the instructions turned out to work better rather than providing the example at the end.
[Context]I am currently working on a task focused on establishing traceability between software requirements and regulatory codes [^3] Note that we use regulatory codes to mean provisions, since the former was used in the literature cleland:2010 . This involves analyzing and mapping requirements to relevant GDPR regulations, ensuring that our software development aligns with regulatory compliance. Below are the main regulatory codes that I want you to remember at first: {The 26 regulatory codes with their descriptions + a 27th code capturing the ‘‘ELSE’’ value indicating no trace link.}
[Examples] Here are five sample traceability examples. I’ve also added my rationale for tracing regulatory codes to the requirements for your reference. {Five example requirements along with their trace links and the rationale behind selecting these links. Requirement: TEXT. trace links: LIST, rational behind choosing these codes: TEXT.}
[Instruction] Find the trace links for a given requirement and provide the rationale behind your choice extended from the examples I provided. Please consider regulatory codes which I have not used in the examples. Pay attention to the roles (AS_ROLE) in the requirement, if there are any. Remember, regulations’ text focus on personal data, but try to consider all types of data, role, or functionalities in a software system. Pay attention to commonsense and indirect relations between requirement and regulations. Aim to include regulations even if they have a low likelihood of being traced, prioritizing recall over precision. Choose at least one regulation for each requirement.
[Output Indicator] List of alphabetical order of regulatory codes (if any) similar to the examples I provided to you. Newline to explain the rational behind the choice(s).
Figure 4: Final Rice_LRT prompt for addressing LRT.
Step 2: LLM Querying
This step applies the prompt designed in Step 1 to instruct the LLM to predict trace links in textual requirements. A prerequisite for using our prompt involves creating a few examples that will demonstrate the LRT task to the LLM. To effectively trigger the reasoning of the LLM, we built five examples by exposing both the labels (i.e., trace links) and the rationale behind selecting these labels. The examples are then integrated into the above LRT prompt template. We applied our prompt by querying the GPT-4o model through the OpenAI API.
Below, we show an example output for a given requirement from KeePass, a password management project [^4] https://keepass.info/extensions/v1/docs/SoftwareRequirementsSpecification-KeePass-1.10.pdf using our prompt (shown in Fig. 4). As mentioned above, we instruct the LLM through our output indicator to generate the rationale for each regulatory code recommended for a given requirement. The rationale is meant to help the analysts decide whether to accept or disregard a trace link. They can examine the reasoning and logic behind these codes generated by the LLM to understand their potential relevance to the requirement and use their knowledge and interpretation of the requirements documents to decide whether the proposed trace links by the LLM apply in their context.
GPT-4o Example Output REQ: If a key file is required to unlock the database, the database doesn’t open unless the key file is present.
Trace links: [ACC, CNF, SEC]
Rationale:
ACC [^5] ACC: Right to access. : The presence of the key file is required to gain access to the database, ensuring proper authentication. CNF [^6] CNF: Ensure confidentiality. : Protecting the database with a key file ensures that sensitive data remains confidential and accessible only to those with the correct authentication file. SEC [^7] SEC: Ensuring security. : Requiring a key file adds a layer of security, preventing unauthorized access even if the password is compromised.
4 Evaluation
In this section, we report on our empirical evaluation.
4.1 Research Questions (RQs)
This paper investigates the following RQs:
RQ1. Which ST model yields the most accurate results for tracing requirements to provisions? As discussed in Section 3, step 2 in $\mathcal{K}$ashif involves selecting the most accurate pre-trained model for the LRT task. Several alternative pre-trained models are publicly available. In RQ1, we examine 38 alternatives reported to work well in the NLP community. The goal of RQ1 is to identify the most accurate ST model for predicting trace links between requirements and provisions.
RQ2. How accurate is $\mathcal{K}$ashif compared to an existing baseline on a standard dataset from the literature? RQ2 aims to assess the value of utilizing ST as enabling technology for addressing the LRT problem compared to a baseline from the existing literature, which we re-implement in this work. The baseline is a classifier that leverages the terminology probability distributions to compute the likelihood that a requirement can be traced to a provision, based on the occurrence of some indicator terms within that provision. The investigation of RQ2 is conducted using the HIPAA dataset.
RQ3. How accurately does $\mathcal{K}$ashif perform on a more complex dataset, spanning multiple requirements types and domains? In RQ3, we test $\mathcal{K}$ashif on four different documents, two shall-requirements and two user stories, covering various domains. These documents are traced to the GDPR privacy requirements. The goal of RQ3 is to investigate the performance of $\mathcal{K}$ashif on a more realistic dataset that captures the complexity of the legal domain.
RQ4. How accurate is Rice_LRT in addressing the LRT task compared to $\mathcal{K}$ashif and existing baselines? Given the recent rise in the use of LLMs, a straightforward alternative for automating tasks such as LRT is to prompt pre-trained LLMs, e.g., GPT-4o, for traceability tasks. RQ4 examines how well our proposed prompt, Rice_LRT, performs compared to the best classifier, $\mathcal{K}$ashif, and existing baselines in the literature.
4.2 Datasets
Table 1: Statistics of the HIPAA dataset cleland:2010 . Rows list the documents in HIPAA, and columns provide their description and the distribution of the trace links across provisions in each document.
ID Description All AC AUD AL EAP PA SED TED TS IC UUI
H1 ClearHealth: EMR System. 44 1 4 1 0 0 1 1 0 2 1
H2 Physician: Electronic Info. Exchange between Clinicians. 147 7 2 0 2 0 0 0 1 3 0
H3 iTrust: Role-based HCT Web app. 184 2 35 1 0 6 0 0 0 0 2
H4 Trial Implementations: National Coordinator for Health IT 100 4 6 0 0 13 0 0 2 4 2
H5 PracticeOne: A Suite of HCT Info. Systems. 34 3 1 0 0 1 0 0 1 1 0
H6 PatientOS: HCT Info. System. 91 1 2 3 1 0 3 1 1 0 1
H7 Lauesen: Sample EMR System. 66 11 0 1 0 5 0 0 0 3 1
H8 WorldVistA: Veteran Administrations EMR. 117 6 2 2 0 4 0 0 0 0 1
H9 Care2x: Hospital Info. System. 44 1 1 1 0 1 1 1 0 0 0
H10 CCHIT: Certification Commission for HCT. 1064 17 33 1 1 12 2 2 2 5 3
Total counts 1891 53 86 10 4 42 7 5 7 18 11
    EMR: Electronic Medical Record. HCT: Healthcare Technology.
We develop our approach and base our initial evaluation on the HIPAA dataset, a publicly available dataset, created and released in 2010 cleland:2010 and reused in 2017 Guo:17 . The dataset was manually created by identifying trace links of requirements against the regulatory statements elicited from the USA government’s Health Insurance Privacy and Portability Act (HIPAA) regulation. The provisions are the following: access control (AC), audit controls (AUD), person or entity authentication (PA), transmission security (TS), unique user identification (UUI), emergency access procedure (EAP), automatic logoff (AL), encryption and decryption (SED), encryption (TED), and integrity controls (IC). HIPAA consists of 10 requirements documents, all shall-requirements, from the healthcare domain. In total, the dataset contains 1,891 requirements, of which 243 have trace links. Table 1 summarizes the different documents (rows) in HIPAA, their description, and the distribution of the trace links across provisions (columns).
Table 2: Test documents used in RQ3 and RQ4
ID Description (S), Domain (D), Number of requirements (N), Number of trace links (T), Type (Y): (1) “Shall” Requirements or (2) User Stories
RD1 S: Keepass is about password management, D: cybersecurity, N: 78, T: 64, Y: 1
RD2 S: WASP is about Functionalities and services provided by the WASP platform, D: digital services, N: 69, T: 73, Y: 1
RD3 S: Datahub is about information on requirements for data publishers, D: digital library systems, N: 66, T: 51, Y: 2
RD4 S: Scrumalliance is about member interactions and data management on professional networking, D: professional development and certification systems, N: 97, T: 93, Y: 2
Table 3: Statistics of the Test documents. Columns list the documents and rows provide their the distribution of the trace links across provisions in each document.
Provision ID Description Title RD1 RD2 RD3 RD4 Total
ACC Right to access 8 22 27 57 114
REC Right to rectification 5 2 3 7 17
RES Right to restriction 0 0 0 0 0
CMP Right to complaint 0 0 0 0 0
ERS Right to erasure 6 2 4 3 15
OBJ Right to object 0 0 0 0 0
PRT Right to portability 1 0 0 0 1
WCON Right to withdraw consent 0 0 0 0 0
CON Consent 0 12 6 10 28
CAT Personal data category 0 16 0 0 16
SCAT Personal data special category 0 0 0 0 0
ORG Personal data origin 0 1 0 0 1
DIR Direct 0 0 0 0 0
PUB Publicly 0 0 0 0 0
TPA Third party 0 0 0 0 0
COK Cookie 0 0 0 0 0
TEC Technical measures 9 0 0 0 9
SEC Ensuring security 25 2 1 1 29
SAS Security assessment 1 0 0 0 1
TRN Personal data transfer 2 0 4 1 7
CHL Children 0 0 0 0 0
TIM Personal data time stored 0 2 0 5 7
DUR Processing duration 4 0 0 5 9
CNF Ensure confidentiality 3 5 0 4 12
BRC Inform breach to data subject 0 4 0 0 4
NTF Data breach notification content 0 5 0 0 5
Total Links 64 73 51 93 281
To address RQ3, we create and curate four documents covering different requirements types and domains. These documents represent a snapshot of a practical scenario that exemplifies the potential complexity of LRT in practice. For each document, we manually identify trace links between software requirements and a list of 26 provisions derived from GDPR and pertinent to software. Building on existing work Amaral:21 ; Amaral:23a , the codes were comprehensively created, in collaboration with a legal expert (non-author), to represent the privacy requirements in GDPR pertinent to software engineering. Table 2 and 3 describe our test documents. Two co-authors of this paper, with more than 10 years of expertise in requirements engineering, manually analyzed the four documents and identified the trace links for all requirements. The test documents originate from four independent sources, each from a different domain, as shown in Table 2.
It is worth mentioning that the test documents are newly created resources that have not yet been publicly released. Although the underlying content may not rely entirely on new concepts, the traceability links between the software projects and the GDPR are new and have not been shared before. Therefore, we can confidently claim that there is no data leakage when addressing this research question using GPT4.
4.3 Baselines
To evaluate our proposed solutions, we compare them against carefully selected baselines. we present the baselines for both $\mathcal{K}$ashif and Rice_LRT, covering their classifier-based and prompt-based variants.
4.3.1 Classifier Baselines
For $\mathcal{K}$ashif, we adopt the following approaches from the literature: (1) well-known information retrieval techniques such as LDA and LSI, (2) a statistical classification approach designed specifically for identifying traceability links in legal documents (we name it B throughout the paper), (3) a more recent static word embedding method, GloVe pennington2014glove , (4) a modern BERT-based model, TraceBERT lin2021traceability , originally developed for detecting trace links between code and issue reports, (5) RoBERTa liu2019roberta , an improved variant of BERT for enhanced language understanding, and (6) a recent technique built upon LLaMa, proposed by Ge et al. ge2025cross , designed to support high-to-low level requirements traceability. This selection ensures a comprehensive comparison across traditional, statistical, static embedding-based, transformer-based, and LLM techniques.
LSI and LDA. Vector Space Model (VSM), LSI, and LDA have been widely adopted in the field of requirements engineering, particularly for addressing the traceability problem guo2025natural , which aims to identify and recover links between related software artifacts such as requirements, design documents, and code. VSM represents each document (e.g., a requirement or a provision) as a vector in a multi-dimensional space, where each dimension corresponds to a term from the overall vocabulary. Similarity between two artifacts is typically computed using the cosine of the angle between their corresponding vectors, capturing the degree of term overlap across all artifacts. LSI extends the Vector Space Model by applying Singular Value Decomposition (SVD) baeza1999modern to the term-document matrix, thereby generating a lower-dimensional latent semantic space. The dimensionality of this subspace is controlled by a manually tuned parameter, commonly referred to as the k value. By capturing the most significant underlying patterns in term usage, LSI implicitly accounts for frequently co-occurring terms, thereby mitigating the vocabulary mismatch problem commonly encountered in information retrieval-based traceability recovery. LDA shares LSI’s goal of uncovering the latent semantic meaning in text, but instead of using SVD, it employs a Bayesian model to identify the underlying structure in word usage across documents.
B. We re-implement as part of this work the baseline B from the literature cleland:2010 ; Guo:17 . B is a probabilistic approach based on occurrences of words in requirements texts and how likely these words are associated with specific provisions. Specifically, B predicts whether a requirement is traced to a provision by identifying keywords (also known as indicator terms) that are present in the requirement. Given an input requirement for which the trace link should be predicted, B requires a training set from which likelihood estimates of indicator terms are computed for the input requirement, representing how likely it is to be relevant to a specific regulation. The training set comprises provisions, software requirements, and the trace links between them. During training, indicator terms are identified and weighted for each provision by parsing the textual requirements traced to these statements. The weights are computed based on factors such as term frequency in related requirements, the fraction of regulation-related requirements containing the term, and the fraction of projects (specific to the HIPAA dataset) involving regulation-related requirements that also contain the term. Given the absence of publicly released implementation for the baseline, we present in this paper a replicated version of B which follows the same procedure described above.
GloVe. We implement a technique based on static word embeddings using GloVe, which generates dense vector representations of words by leveraging global co-occurrence statistics from a large corpus, thereby capturing both semantic and syntactic relationships between words. In our work, we use GloVe as a trainable embedding layer and as the input layer in a simple neural network architecture (with a single layer to predict the score). In this context, the inputs consist of a requirement and a candidate provision, and the output is a probability score computed using a sigmoid activation function. The model is subsequently fine-tuned on the training dataset, allowing it to adapt the pre-trained word embeddings to the specific characteristics of the traceability task.
TraceBERT. we utilize the publicly available implementation of TraceBERT, a transformer-based model that, like our approach, is built on a transformer architecture. In this paper, the authors propose three distinct BERT-based architectures (TWIN, Single, and Siamese) for fine-tuning the pre-trained BERT model specifically for the traceability link prediction task. Each architecture represents a different way to structure the input and optimize the model for semantic similarity between software artifacts. Based on reported evaluation results, the Single-TraceBERT architecture outperforms the other variations, and we adopt it as one of the baseline models in this work. Note that, unlike the original TraceBERT paper, which includes intermediate training on CodeSearchNet husain2019codesearchnet , we skipped this step as it is not relevant to legal domain traceability as CodeSearchNet contains programming language artifacts which is beneficial for source code and issue traceability, but differs fundamentally—in both terminology and syntactic structure—from establishing traceability between two NL artifacts in our case. Therefore, we directly fine-tuned the authors’ architecture on our dataset.
[…]
LLaMa. Finally, we utilize the publicly available implementation of a recent approach by Ge et al. ge2025cross , which primarily focused on establishing traceability between low- and high-level requirements using LLaMa2 touvron2023llama . In this work, they use data augmentation to fine-tune LLaMa2 at different scales (1.1B, 7B, and 13B) and demonstrate that LLaMa2-7B achieves the best performance. They demonstrate that the fine-tuned LLaMa outperforms both IR-based and ML-based approaches on six datasets, using a 90/10 train–test split. In our study, we evaluate how well this approach performs on the requirement–provision task under a more complex evaluation setting.
4.3.2 Prompt Baselines
For Rice_LRT, we incorporate four baselines, P1, P2, P3_1, and P3_2, from the literature hey2025requirements ; ronanki2024requirements , which focus on identifying links between high-level and low-level requirements. This comparison enables us to assess the effectiveness of our context-rich prompt, Rice_LRT, relative to recent prompt-based techniques reported in the literature for traceability.
P1. As part of this work, we reimplemented an approach based on Retrieval-Augmented Generation (RAG) initially proposed for inter-requirement traceability hey2025requirements . This RAG-based approach consists of two main steps: retrieval and prompting. In the retrieval step, semantic similarity is used to select the top-k target artifacts that are most similar to a given source artifact. For the prompting step, the authors employed two types of prompts: KISS, a simple zero-shot prompt, and CoT, a chain-of-thought prompt. Based on their results, CoT prompting demonstrated superior performance. Therefore, in this work, we focus on CoT and introduce a minor prompt modification to better suit our LRT task. In our context, the source artifacts are the requirements, and the target artifacts are the provisions, respectively. The prompt used is:
Below are artifacts from a software system requirement and the GDPR. Is there a traceability link between (1) and (2)? Give your reasoning and then answer with ’yes’ or ’no’ enclosed in <<trace>> <</trace>>.
(1) Requirement: '''{source_content}'''
(2) Regulation: '''{target_content}'''
P2. We adopt the prompt template proposed by Ronanki et al. ronanki2024requirements . Their study presents five templates for identifying traceability between requirements, two of which require human involvement, whereas this paper does not. From the remaining three templates, we select the one that achieved the highest F2 score. We additionally introduce minor modifications to tailor the prompt to the specific needs of our LRT task. The prompt used is as follows:
Act as a requirements engineering domain expert and list the IDs of the GDPR regulations that are dependent on the following requirement:
Requirement: '''{content}'''
List of Regulations: '''{List}'''
P3. We adopt two prompt templates used by Ge et al. ge2025cross as part of their investigation into high- and low-level requirement traceability. Their study evaluated four zero-shot prompt variations, ranging from a simple prompt to a more detailed CoT formulation. In order to have a variety of prompts in our baselines, we employ the prompt with the best results in their work, we call it P3_1 throughout the paper, along with the CoT-based prompt as it includes more detailed instructions compared to P1 and P2, we call it P3_2. We modify the prompt to better fit the LRT task. The prompts we used are as follows:
P3-1: Consider the following scenario where the Requirement text represent requirements of the software system and Regulation represent a GDPR regulation. Let’s think step by step: 1. Identify the key elements of the requirement; 2. Match these elements with the regulation; 3. Highlight any gaps; 4. Infer implicit traceability; Finally, answer with Yes or No if there is a traceability relationship, and provide a brief rationale in about 10 words.
Requirement: '''{content}'''
Regulation: '''{content}'''
P3-2: Consider the following scenario where the Requirement text represent requirements of the software system and Regulation represent a GDPR regulation. Answer with Yes or No: Does the regulation trace back to the requirement?
Requirement: '''{content}'''
Regulation: '''{content}'''
4.4 Implementation
$\mathcal{K}$ashif. We implement $\mathcal{K}$ashif in Python 3.8. For preprocessing the text, we use the NLTK toolkit (v 3.8.1). We access the ST pre-trained models through the Hugging Face Transformers library (4.44.0). For fine-tuning, we use the Sentence-Transformers library (2.6.1). We use the same library for computing cosine similarity. Our experiments were performed on an RTX 6000 GPU with 24 GB of RAM. For fine-tuning, we relied on the ST29 base model. We performed grid search to tune the hyperparameters, considering epochs=[5, 10, 15], batch size=[8, 16, 32], optimizer=[AdamW, RMSprop], learning rate=[2e-5, 2e-3, 2e-2], weight decay=[2e-3, 2e-2, 2e-1], and warmup percentage=[5e-2, 5e-1]. The best-performing hyperparameters were epochs=5, batch size=8, optimizer=AdamW, linear learning rate=2e-5, weight decay=2e-1, and warmup percentage=5e-2. Checkpoints were saved at the end of each epoch, and the final model was selected based on the lowest observed loss.
RICE_LRT. We implement Rice_LRT in Python 3.8. using the OpenAI (1.97.1) API with the following settings: a temperature of 0, a max-token of 2,000, a frequency penalty of 0, a presence penalty of 0, top_p = 1, and a random seed of 16.
B. We also implement B in Python 3.8. We have used the scikit-learn library (1.7.0) to implement the probabilistic functions.
LSI and LDA. For LSI and LDA, we utilized the scikit-learn library (v1.7.0), setting the number of components to 50 within a search space of 5, 10, 25, and 50 components. For LDA the Document–Topic Dirichlet Prior ($\alpha$) and Topic–Word Dirichlet Prior ($\beta$) had the values of 2e-2 within a search space of 1e-2, 2e-2, and 1e-1. For both models, we employed TF-IDF as the feature weighting scheme using scikit-learn library (v1.7.0). Our experiments showed that the LSI model achieved optimal performance when stopwords and punctuation were removed. In contrast, the LDA model performed best when stopwords and punctuation were removed and lemmatization was applied using NLTK toolkit (v3.8.1). We did not apply minimum cutoff thresholds for the vocabulary, as filtering out low-frequency terms could risk omitting important information relevant to some regulatory provisions. For the maximum document frequency cutoff, we experimented with values of 0, 1e-1, and 1. We observed that as the cutoff value increased, the model’s performance decreased. Therefore, we decided not to apply a maximum cutoff in our configuration.
GloVe. For GloVe, we used it as input to a single-layer neural network with a sigmoid activation function to predict the similarity score implemented in TensorFlow (2.18.0). We performed grid search to tune the hyperparameters, considering epochs=[5, 10, 15], output dimension=[150, 300], max_length=[50, 100, 150]. The best-performing hyperparameters were using epochs=10, output dimension=300, and max_length=50.
TraceBERT. For TraceBERT, we used the publicly available code from data preparation for tuning the model and retained the fine-tuning settings in its shared implementation.
RoBERTa. For RoBERTa, the entire fine-tuning setup is implemented using the Hugging Face Transformers library. We performed grid search to tune the hyperparameter, considering using epochs=[5, 10, 15], batch size=[8, 16, 32], optimizer=[AdamW, RMSprop], learning rate=[1e-5, 2e-5, 1e-3, 2e-3], and weight decay=[0, 1e-5, 1e-3, 1e-2, 1e-1]. The best-performing hyperparameters were epochs=5, batch size=32, optimizer=AdamW, learning rate=2e-5, and weight decay=1e-2.
LLaMa. For LLaMa, we used the publicly available implementation presented by Ge et al ge2025cross . We relied on their released code to generate the training set following the procedure outlined in their paper. Since using the same hyperparameters resulted in a single label being produced for all pairs, we performed another round of hyper-parameter tuning considering learning rate=[2e-8, 2e-7, 2e-5, 2e-5, 2e-4, 2e-3, 2e-2] and weight decay=[0, 1e-5, 1e-3, 1e-2, 1e-1], which the best-performing hyperparameters were learning rate=2e-4 and weight decay=1e-5.
P1. For P1, we re-implemented the approach following the settings provided by the authors. The prompt was executed on GPT-4o with a temperature of 0 and a random seed of 133742243. For the RAG step, we used the same embeddings as the authors proposed, specifically the OpenAI embedding text-embedding-3-large.
P2. For P2, we re-implemented the approach following the settings provided by the authors. While the original paper used GPT-3.5-turbo, our comparative experiments revealed that GPT-4o yielded better outcomes. Therefore, the results for P2 presented in this paper are based the GPT-4o model. Moreover, the choice of random seed and temperature followed the best-performing settings reported in their paper.
P3. For P3_1 and P3_2, we re-implemented the approach using the settings provided by the authors over GPT-4o.
4.5 Pre-trained Model Selection (RQ1)
Methodology. We shortlist the ST models for investigation in our work based on the NLP leaderboard, which reports the 38 most accurate pre-trained models [^8] https://www.sbert.net/docs/pretrained_models.html. These models have been extensively evaluated for their ability to generate sentence embeddings (i.e., capturing the semantics of the whole text) and their performance in semantic search (i.e., finding relevant answers to a given query). Both tasks closely align with our objectives. To identify trace links, we apply the pre-trained models in a zero-shot setting as follows. We let each model compute the similarity matrix equivalent to the output of step 5 in our approach (see Fig. 2). We then predict a trace link if the similarity value exceeds a predefined threshold. Since zero-shot does not require training, we simply run the pre-trained model on the entire HIPAA dataset.
Table 4: AUC of ST models for LRT on HIPAA (RQ1)
KK1 Model2 Name1 AUC1 $K^{\dagger}$1
1 ST1 all-mpnet-base-v2 0.744 16
2 ST2 gtr-t5-xxl 0.725 21
3 ST3 gtr-t5-xl 0.789 6
4 ST4 sentence-t5-xxl 0.720 22
5 ST5 gtr-t5-large 0.743 17
6 ST6 all-mpnet-base-v1 0.712 25
7 ST7 multi-qa-mpnet-base-dot-v1 0.688 27
8 ST8 multi-qa-mpnet-base-cos-v1 0.603 34
9 ST9 all-roberta-large-v1 0.601 35
10 ST10 sentence-t5-xl 0.769 10
11 ST11 all-distilroberta-v1 0.719 23
12 ST12 all-MiniLM-L12-v1 0.729 19
13 ST13 all-MiniLM-L12-v2 0.747 15
14 ST14 multi-qa-distilbert-dot-v1 0.563 36
15 ST15 multi-qa-distilbert-cos-v1 0.640 33
16 ST16 gtr-t5-base 0.770 9
17 ST17 sentence-t5-large 0.748 14
18 ST18 all-MiniLM-L6-v2 0.761 11
19 ST19 multi-qa-MiniLM-L6-cos-v1 0.670 29
20 ST20 all-MiniLM-L6-v1 0.749 13
21 ST21 paraphrase-mpnet-base-v2 0.850 2
22 ST22 msmarco-bert-base-dot-v5 0.644 32
23 ST23 multi-qa-MiniLM-L6-dot-v1 0.715 24
24 ST24 sentence-t5-base 0.726 20
25 ST25 msmarco-distilbert-base-tas-b 0.701 26
26 ST26 msmarco-distilbert-dot-v5 0.685 28
27 ST27 paraphrase-distilroberta-base-v2 0.801 4
28 ST28 paraphrase-MiniLM-L12-v2 0.794 5
29 ST29 paraphrase-multilingual-mpnet-base-v2 0.859 1
30 ST30 paraphrase-TinyBERT-L6-v2 0.787 7
31 ST31 paraphrase-MiniLM-L6-v2 0.770 8
32 ST32 paraphrase-albert-small-v2 0.737 18
33 ST33 paraphrase-multilingual-MiniLM-L12-v2 0.811 3
34 ST34 paraphrase-MiniLM-L3-v2 0.757 12
35 ST35 distiluse-base-multilingual-cased-v1 0.349 37
36 ST36 distiluse-base-multilingual-cased-v2 0.341 38
37 ST37 average_word_embeddings_komninos 0.647 31
38 ST38 average_word_embeddings_glove.6B.300d 0.636 30
  • 1
    $K$: The average performance ranking of the models, as reported in the NLP community. $K^{\dagger}$: The ranking of the models based on AUC values computed on HIPAA ($K=1$ indicates the highest AUC).
  • 2
    ST1ST38 correspond to the models reported at this link (sorted by average accuracy in descending order): https://www.sbert.net/docs/pretrained_models.html.
Evaluation Metrics. To better assess the performance irrespective of the selected threshold, we compute the Area Under the Curve (AUC) for the receiver operating characteristic (ROC) across different threshold values, ranging from $0.1$ to $0.9$. The ROC curve captures the trade-off between the true positive rate (TPR) and the false positive rate (FPR). TPR is the proportion of positives correctly identified as such (i.e., the percentage of trace links correctly identified for a given threshold). FPR is the proportion of negatives incorrectly identified as positives (i.e., the percentage of trace links wrongly identified as not trace links). The AUC of the ROC curve (computed as micro-average over all the provisions to avoid the dominance of some provisions) provides a single aggregate performance measure across all possible thresholds and, hence, is a suitable evaluation metric to compare the ST models. We posit that the model with the highest AUC value demonstrates the best overall performance in identifying trace links in a zero-shot setting, as a higher AUC value indicates a better balance between correctly identifying true trace links (high TPR) and minimizing the identification of false links (low FPR).
Results. Table 4 presents the AUC values of the ST pre-trained models on the HIPAA dataset and also reports $K$, indicating the ranking of the models in the NLP community based on their accuracy Reimers:19 , as well as $K^{\dagger}$, indicating the ranking based on AUC achieved on HIPAA.
The best-performing model on HIPAA is ST29 ($K^{\dagger}=1$), with an AUC value of 0.859. The next best performing model is ST21 with an AUC value of 0.850. The difference between these two AUC values is only marginal. A possible explanation is that ST29 uses ST21 as its base model. ST29 has been, however, trained on more (multi-lingual) data.
Additionally, we observe a discrepancy in the performance of the models on the HIPAA dataset compared to that reported by the NLP community. The best NLP model, ST1, does not perform well on HIPAA, ranked 16. This observation indicates that well-performing models in NLP are not necessarily as effective for RE-specific problems. We acknowledge that zero-shot performance does not always predict fine-tuning effectiveness, particularly across different model architectures. To address this concern, we fine-tuned the top five models based on their zero-shot AUC scores to evaluate their performance more comprehensively. However, our experiments show that ST29 still consistently outperforms the other models after fine-tuning. While resource constraints prevented us from fine-tuning all possible models, this targeted evaluation provides evidence that ST29 is the most effective choice among the top-performing candidates.
The answer to RQ1 is that ST29 is the best-performing pre-trained model for LRT (corresponding to paraphrase-multilingual-mpnet-base-v2).
4.6 Accuracy on Benchmark Dataset (RQ2)
Methodology. We compare the four variants of $\mathcal{K}$ashif (explained in Section 3) against LSI, LDA, B from the literature cleland:2010 ; Guo:17 , GloVe, TraceBERT, RoBERTa, and LLaMa. We answer RQ2 on the benchmark dataset, HIPAA. Since HIPAA contains 10 requirements documents, we apply the leave-one-out (LOO) evaluation method, where $\mathcal{K}$ashif and baselines are tested on a left-out document each time and trained (or fine-tuned) on the remaining documents to emulate realistic settings. However, to ensure a reasonable balance between the training and test sets, we exclude one document (CCHIT, labeled H10 in Table 1) from the LOO process, as it contains 1,064 requirements, thereby including more than half the dataset.
Evaluation Metrics. We evaluate the four variations of $\mathcal{K}$ashif and the baselines using precision (P), measuring how many trace links identified by the approach are correct; recall (R), measuring how many trace links in our ground truth are correctly identified by the approach; and F2 score, the harmonic mean of precision and recall, with a greater emphasis on recall. We report the mean and standard deviation across the nine documents. To ease comparisons of the models’ rankings, we also include Mean Average Precision (MAP) over requirements that have at least one trace link to a provision, which measures how well a model ranks relevant items higher than irrelevant ones. We also report the AUC value for the overall performance of the classifiers. Moreover, we report the AUC value for each classifier to better evaluate their performance in distinguishing positive and negative pairs across thresholds. Note that the MAP score is computed only for requirements containing at least one true positive, since requirements without any trace links provide no informative contribution to the MAP. In contrast, AUC is applied to all requirements as it allows for a more complete assessment of each classifier’s capability to distinguish relevant from non-relevant pairs, regardless of the presence or absence of TPs for a particular requirement.
Table 5: Accuracy of $\mathcal{K}$ashif and baselines on HIPAA (RQ2)
TP FP FN P R F2 MAP
$\mathcal{K}$ashif $_{\text{constant}}$ 111 $\pm 12$ 114 $\pm 8$ 54 $\pm 4$ 49.3 $\pm 13$ 67.3 $\pm 18$ 62.7 $\pm 11$ 81.4
$\mathcal{K}$ashif $_{\text{dynamic}}$ 122 $\pm 12$ 441 $\pm 46$ 43 $\pm 4$ 21.7 $\pm 18$ 73.9 $\pm 21$ 49.0 $\pm 13$
$\mathcal{K}$ashif Δ 132 $\pm 12$ 1531 $\pm 81$ 33 $\pm 2$ 7.9 $\pm 3$ 80.0 $\pm 12$ 28.3 $\pm 4$
$\mathcal{K}$ashif $_{\text{tuned}}$ 94 $\pm 11$ 44 $\pm 4$ 71 $\pm 3$ 68.1 $\pm 19$ 56.9 $\pm 17$ 58.5 $\pm 12$
LDA 66 $\pm$11 1571 $\pm$253 99 $\pm$6 4.1 $\pm$14 40.0 $\pm$24 14.4 $\pm$3 37.8
LSI 19 $\pm$2 202 $\pm$31 146 $\pm$11 8.5 $\pm$18 11.5 $\pm$17 10.7 $\pm$13 53.5
B 24 $\pm$2 16 $\pm$1 141 $\pm$10 60.0 $\pm$20 14.5 $\pm$9 17.1 $\pm$11 78.8
Glove 50 $\pm$11 428 $\pm$55 115 $\pm$7 10.5 $\pm$8 30.3 $\pm$25 22.0 $\pm$11 48.1
TraceBERT 15 $\pm$3 97 $\pm$20 147 $\pm$12 13.4 $\pm$6 9.3 $\pm$25 9.8 $\pm$8 62.7
RoBERTa 129 $\pm$7 784 $\pm$39 36 $\pm$6 14.1 $\pm$1 78.1 $\pm$12 41.0 $\pm$1 55.9
LLaMa2-7b-w/o* 33 $\pm$6 1366 $\pm$229 129 $\pm$11 2.4 $\pm$1 20.3 $\pm$31 8.1 $\pm$5 28.3
LLaMa2-7b-tuned 111 $\pm$12 1713 $\pm$102 51 $\pm$3 6.1 $\pm$2 67.3 $\pm$17 22.5 $\pm$7 41.4
  • 1
    *: w/o: without fine-tuning, rather applied in zero-shot setting.
Results. $\mathcal{K}$ashif $_{\text{tuned}}$ and the baseline models (LDA, LSI, B, and GloVe) require a decision threshold, which we tune using the same strategy explained in Section 3. In contrast, RoBERTa relies on the softmax function inherent to its architecture for final class-label assignment, and tuning the threshold over the output logits for label 1 resulted in a lower F2 score. Similarly, the publicly available implementations of TraceBERT and LLaMA output both class probabilities and a final label, and using the label as the final output yielded a higher F2 score. Therefore, these models do not require separate threshold selection. To compute the MAP value for RoBERTa, we use the output logits (scores) corresponding to the [CLS] token for label 1; for TraceBERT and LLaMa, we use the probability outputs.
Table 5 lists, for each approach, the total number of TPs, FPs, FNs, and TNs, and further reports the mean and standard deviation of precision, recall, F2, and MAP. Moreover, Figure 5(a) and Figure 5(b) present the ROC-AUC curve and the average F2 score along 200 evenly spaced threshold values between 0 and 1 for each classifier, respectively.
As visible from Table 5, $\mathcal{K}$ashif $_{\text{constant}}$ has the highest F2 score compared to the baselines. $\mathcal{K}$ashif also has the highest MAP score (81.4%), indicating the best ranking performance: in top-k results, it has the most TPs among baselines. Moreover, $\mathcal{K}$ashif has the highest AUC (0.93), indicating greater ability to distinguish between positive and negative pairs. $\mathcal{K}$ashif $_{\text{tuned}}$ outperforms all variants of $\mathcal{K}$ashif in terms of precision, achieving an average of 68.1%. Based on our observations of the tuned thresholds, we found that most threshold values are greater than 0.5, which explains the higher precision compared to $\mathcal{K}$ashif $_{\text{constant}}$. For the H1 to H9 projects, the corresponding threshold values are 0.91, 0.5, 0.79, 0.46, 0.82, 0.81, 0.78, 0.73, and 0.58, respectively.
The IR-based methods, LSI and LDA, exhibit notably different behaviors. LDA achieves an F2 score of only 14.4% and a MAP of 37.8%, indicating its failure to assign higher ranks to true-positive pairs. Furthermore, its AUC of 0.51 suggests that LDA performs no better than random and offers little discriminatory power. In contrast, LSI attains an F2 score of 10.7%, which is 4 pp lower than LDA. However, it outperforms LDA in both MAP (53.5%) and AUC (0.71). This suggests that LSI has stronger ranking capability. Additionally, Figure 5(b) shows that LSI’s maximum achievable F2 score is close to 20%, whereas LDA’s peak F2 score is around 14%. Therefore, the threshold-tuning process did not yield the best LSI performance.
B, the ML-based technique tailored for the LRT task, exhibits relatively low retrieval performance compared to the other classifiers. Nevertheless, it achieves the second-highest MAP score (78.8%), following $\mathcal{K}$ashif, indicating a strong ability to rank positive pairs higher than negative ones. We recall that B is a classifier that primarily uses a probabilistic approach based on word co-occurrence in requirement texts to predict whether each requirement should be traced to a particular provision. This indicates that B has greater discriminative ability between positive and negative pairs than IR-based techniques, GloVe, TraceBERT, and LLaMa. Moreover, Figure 5(b) shows that B’s maximum achievable F2 score is around 38%, indicating that threshold tuning was unable to identify an effective decision boundary for model B.
The GloVe classifier achieves an F2 score of 22%, a MAP of 48.1%, and an AUC of 0.73. These results indicate that GloVe behaves similarly to LSI. Unlike some other models, threshold tuning for GloVe appears effective, as shown in Figure 5(b). Its maximum attainable F2 score is approximately 22–23%, and the selected tuned thresholds achieve an F2 score of 22%.
TraceBERT, however, performs poorly on the LRT task, with results comparable to LDA. It achieves an F2 score of 8.7% and an AUC of 0.67, results that are only marginally better than random. However, its relatively high MAP of 62.7% indicates that it learned the relative similarity between pairs. We observed that TraceBERT produces nearly identical similarity values for most requirement-provision pairs, around 0.5. As a result, its threshold-based classification is substantially hindered. Moreover, in Figure 5(b), TraceBERT’s F2 score starts to change at a threshold value around 0.5. Recall that TraceBERT was originally designed for issue–code traceability and incorporates an intermediate training stage on a large external dataset. Therefore, its inability to achieve high performance in our setting is not unexpected.
RoBERTa achieves an F2 score of 41%, a MAP of 55.9%, and an AUC of 0.88, the second-highest AUC among all classifiers. Its high AUC value indicates that RoBERTa effectively separates positive and negative pairs. However, its ranking quality shows greater variability than $\mathcal{K}$ashif and B, as reflected in its MAP, which is fourth-highest overall. RoBERTa exhibits an almost constant F2 score across different threshold values, as shown in Figure 5(b). The maximum attainable F2 score for RoBERTa is around 41–43%, suggesting that the softmax-based output layer is already calibrated to produce its best possible performance within this range. Among the baselines, RoBERTa achieves the highest F2 score. RoBERTa also has the highest recall compared to all models and $\mathcal{K}$ashif variations, except for $\mathcal{K}$ashif Δ. However, compared to $\mathcal{K}$ashif constant, it has a lower recall, which shows weaker effectiveness in distinguishing relevant links by retrieving many FPs (650 FP links more than $\mathcal{K}$ashif constant).
Regarding the use of LLaMa as a classifier, we observe that LLaMa-w/o performs poorly with the lowest F2 score. After fine-tuning, however, the model’s performance improves significantly, demonstrating that it has learned task-specific cues. LLaMa-w/o has the AUC value of 0.50, which shows a random behavior and lacks discrimination on the LRT task. However, LLaMa-tuned achieves an F2 score of 22.5%, an AUC of 0.80, and a MAP of 41.4%, indicating that while the model can distinguish between positive and negative pairs (as reflected in its relatively high AUC), its overall ranking and threshold-based performance remain limited. The low MAP value suggests that LLaMa-tuned struggles to consistently assign high similarity scores to all true-positive pairs, capturing only a subset of links while missing many others. Moreover, in Figure 5(b), the highest attainable F2 score for LLaMa-tuned is above 25% and lower than 30%.
According to Figure 5(b), threshold tuning plays a crucial role in achieving optimal F2 scores for classifiers that rely on a threshold for final label assignment. For instance, B reaches an F2 score of only 17.1% when using a threshold tuned on the training set, even though its maximum attainable F2 score is approximately 38%. This behavior can also be observed for LSI. In contrast, models such as RoBERTa, which rely on softmax-based probability outputs for label assignment, already operate near their maximum F2 scores, as evidenced in Table 5 and Figure 5(b). On the other hand, for $\mathcal{K}$ashif, the optimal F2 scores are primarily concentrated within the 0.4–0.7 threshold range, which is due to its fine-tuning process described in Section 3. Not only does $\mathcal{K}$ashif achieve the highest F2 score compared to all classifiers, but it also offers the practical advantages of being stable around a 0.5 threshold range and thus reduces the need for threshold tuning.
(a) ROC-AUC plot for all classifiers.
(b) F2 scores over different threshold values.
Figure 5: Evaluating classifier performance across different thresholds.
Comparing the four variants of $\mathcal{K}$ashif, our results show that $\mathcal{K}$ashif $_{\text{constant}}$ is the best performing variant in terms of F2, achieving an average score of 62.7%. This score provides a gain of 4.2 pp over $\mathcal{K}$ashif $_{\text{tuned}}$, 13.7 pp over $\mathcal{K}$ashif dynamic, and 34.4 pp over $\mathcal{K}$ashif Δ. In terms of recall, however, $\mathcal{K}$ashif Δ achieves the best value of 80%, 12.7% more than $\mathcal{K}$ashif $_{\text{constant}}$. This can be explained by the threshold adjustment method for $\mathcal{K}$ashif Δ. Recall from Section 3 that to determine the threshold above which a trace link is predicted, we look at the largest gap in similarity values between the requirement and the provisions. Once determined, $\mathcal{K}$ashif Δ will always predict at least one trace link for each requirement corresponding to the provision with the highest similarity value that exceeds this gap. Such a method and recall value can indeed be useful when building recommendation systems. However, they come at the cost of introducing more FPs (as evidenced by the low precision), requiring the human analyst to filter them out. Consequently, we select $\mathcal{K}$ashif $_{\text{constant}}$ as the best performing model for LRT.
To understand the sources of errors produced by $\mathcal{K}$ashif $_{\text{constant}}$, we analyzed the results per document and provision. The results are listed in Table 6. Our analysis reveals the following causes of errors:
  • $\bullet$
    Computing low similarity scores for existing trace links. A majority of FNs (36/54 = 66.7%) are due to computing low similarity scores between the requirement and the corresponding traced provisions. These low scores do not exceed the threshold, thus leading to FNs.
  • $\bullet$
    Computing high similarity scores when there are no trace links. A majority of FPs (96/113 = 84.9%) are due to falsely predicting a trace link for those requirements that have no trace links in our ground truth. This case suggests that a binary classifier could help in reducing FPs by predicting whether a requirement should have a trace link or not in the first place. We have conducted several experiments around this hypothesis. While we observed fewer FPs with a binary classifier, the overall improvement was not significant; hence, we do not report it in this paper.
  • $\bullet$
    Predicting wrong provisions as trace links. The remaining FPs and FNs are caused by predicting provisions other than those identified in the ground truth.
Table 6: Results of $\mathcal{K}$ashif ($\theta>0.5$) per document and provision
[t] AC AUD AL SED EAP TP FP FN TP FP FN TP FP FN TP FP FN TP FP FN H1 1 0 0 3 12 1 1 0 0 1 1 0 0 0 0 H2 1 2 6 1 0 1 0 0 0 0 0 0 1 0 1 H3 2 9 0 35 4 0 1 0 0 0 8 0 0 3 0 H4 1 2 3 6 0 0 0 0 0 0 0 0 0 0 0 H5 2 1 1 1 1 0 0 0 0 0 0 0 0 0 0 H6 1 10 0 2 0 0 3 1 0 3 0 0 1 0 0 H7 9 5 2 0 2 0 1 0 0 0 0 0 0 0 0 H8 4 10 2 2 1 0 1 0 1 0 0 0 0 0 0 H9 0 1 1 1 0 0 1 0 0 0 0 1 0 0 0 $\sum$ 21 40 15 51 20 2 8 1 1 4 9 1 2 3 1 TED IC PA TS UUI TP FP FN TP FP FN TP FP FN TP FP FN TP FP FN H1 0 1 1 0 1 2 0 0 0 0 0 0 0 0 1 H2 0 0 0 2 0 1 0 0 0 0 6 1 0 0 0 H3 0 0 0 0 1 0 5 1 1 0 1 0 2 3 0 H4 0 0 0 3 3 1 8 3 5 0 0 2 0 3 2 H5 0 0 0 0 0 1 0 0 1 1 3 0 0 1 0 H6 1 2 0 0 0 0 0 4 0 0 0 1 0 0 1 H7 0 0 0 0 0 3 2 1 3 0 0 0 0 0 1 H8 0 0 0 0 2 0 0 1 4 0 0 0 0 3 1 H9 0 0 1 0 0 0 1 0 1 0 0 0 0 0 0 $\sum$ 1 3 2 5 7 8 16 10 14 1 10 4 2 10 6
    See Table 1 for the names of the documents
The answer to RQ2 is that $\mathcal{K}$ashif yields the best accuracy on HIPAA when we apply a constant threshold value of 0.5. Specifically, $\mathcal{K}$ashif achieves an F2 score of $\approx$63%. Compared to the best existing baseline from the literature, $\mathcal{K}$ashif has a gain of about 34 pp in F2 score. Moreover, $\mathcal{K}$ashif achieves the highest MAP score of 0.81 and AUC value of 0.93 compared to the baseline classifiers.
4.7 Effectiveness of Classification (RQ3)
Methodology. In real-life scenarios, dealing with LRT involves navigating through many provisions, usually significantly more than 10, as in the simple HIPAA case. This inherent complexity is notable with the 26 provisions pertinent to software in the GDPR. Using the test documents described in Section 4.2, we evaluate and compare two models, namely ST29—the best pre-trained ST model selected in RQ1 and $\mathcal{K}$ashif $_{\text{constant}}$—the best $\mathcal{K}$ashif variant fine-tuned on HIPAA identified in RQ2. Note that we opted not to fine-tune $\mathcal{K}$ashif again on the new documents for three reasons. First, the documents are small and thus inadequate for meaningful training (or fine-tuning). Second, we aim to challenge existing solutions with a more realistic scenario and apply them to new, unseen documents. Finally, $\mathcal{K}$ashif is a similarity-based solution which has been exposed to both the LRT task as well as the regulatory domain (terminology) in the first fine-tuning on HIPAA. Therefore, another fine-tuning is less likely to have any additional value.
Evaluation Metrics. To evaluate the models’ effectiveness, we report results at the requirements and trace-link levels.
At the requirements level, we report (i) the number of requirements where the recommendations made by the model were exactly the same as our ground truth (exact match); (ii) the number of requirements that were a partial match to the ground truth, i.e., the requirements where the model recommended the same regulatory codes as in the ground truth along with additional recommendations (FP); (iii) the number of incorrect matches, i.e., all the other requirements that are not exact or partial matches. Following this, we compute the success rate as the ratio of requirements for which the approach predicts correct trace links (including both partial and exact matches) to the total number of requirements.
We also want to evaluate, on average, how well the model performs for a single requirement. Thus, we report two additional metrics: (iv) Macro Recall (MR), presented in Equation 1, to assess how well the predicted provisions are actually traced to the requirement; (v) Cost, presented in Equation 2, to capture the effort required for an analyst to check the predicted results and filter out the FPs.
$MR=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\frac{TP_{r_{i}}}{TP_{r_{i}}+FN_{r_{i}}}$ (1)
$Cost=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\frac{TP_{r_{i}}+FP_{r_{i}}}{N_{p}}$ (2)
$N_{r}$ denotes the number of requirements in the dataset, $N_{p}$ the total number of provisions in the GDPR, $TP_{r_{i}}$ and $FP_{r_{i}}$ the number of true positives and false positives predicted as a trace link for requirement $i$. A higher MR value indicates a more complete set of links, with the maximum possible value being 1 (100%). If a requirement has no trace links and the predicted set is also empty, the recall is defined as 1. A lower Cost corresponds to more efficient predictions, thereby reducing analyst effort.
At the trace link level, we report the total number of actual trace links, true positives (TP) and false positives (FP), recall (R), precision (P), and F2 score.
Table 7: Accuracy of ST29 and $\mathcal{K}$ashif on the test documents (RQ3).
[t] Trace Link Level ST29 $\mathcal{K}$ashif N T TP FP R P F2 TP FP R P F2 RD1 73 57 0 1 0.0 0.0 nan 10 95 17.5 9.5 15 RD2 64 65 1 3 0.2 25.0 24.9 11 72 16.9 13.2 16 RD3 61 43 0 15 0.0 0.0 nan 7 69 16.3 9.2 14.1 RD4 92 86 2 1 0.1 66.7 12.5 8 94 9.3 7.8 8.9 Average - - - - 0.1 22.9 9.3 - - 15 9.9 13.5
Requirement Level
ST29 $\mathcal{K}$ashif
N T EM PM SR MR C EM PM SR MR C
RD1 73 57 32 1 45.2 0 0.09 19 16 47.9 10.2 5.1
RD2 64 65 30 0 46.9 0.9 0.02 29 5 54.7 15.3 4.6
RD3 61 43 23 4 44.3 0 0.1 13 16 47.5 18.9 5.9
RD4 92 86 20 0 21.7 4.2 0.03 14 10 26.1 7.4 3.0
Average - - - - 39.5 1.3 0.06 - - 44.1 12.9 4.7
  • T: Predicted trace links, EM: Exact Match, PM: Partial Match, SR: Success Rate, MR: Macro Recall, C: Cost.
Results. Table 7 shows the results for each approach across the test documents, both at the trace link level and at the requirement level. The table shows the number of requirements in each test document [^9] Note that we leave out five requirements from each document to enable fair comparison with the Rice_LRT approach presented in RQ4, which relies on few-shot learning., the number of predicted trace links (T), TPs, FPs, R, P, and F2. It also includes the number of requirements with exact matches, the number with partial matches, and the success rate. From the table, we observe that ST29 performs worse than $\mathcal{K}$ashif in terms of success rate. While ST29 achieves a higher number of exact matches, it produces fewer partial matches compared to $\mathcal{K}$ashif. Note that, for both techniques, exact matches often represent “no trace link”, i.e., not predicting any trace link for requirements that had no trace links according to our ground truth. This highlights ST29’s failure to produce a non-empty set of trace links, resulting in a lower success rate. Therefore, allowing more partial matches to achieve a higher success rate, even at the cost of fewer exact matches, ensures broader coverage of correct links. Indeed, the effort required for a requirements analyst to filter out FPs is relatively low compared to the consequences of a requirement not being traced to a complete set of predicted provisions. Therefore, achieving a higher number of partial matches, despite FPs, can be more beneficial than achieving a higher number of exact matches for supporting requirements-level analysis. Our results indicate that the ST pre-trained model (ST29) failed to automatically predict trace links in most cases, suggesting that it neither understood the LRT task nor the application domain. Moreover, ST29 has the Cost of 0.06% and a MR of 1.3% per requirement. This means that for a requirement, on average, the model retrieves 0.06% of the provisions while ensuring 1.3% of the actual trace links are presented in those retrieved links. $\mathcal{K}$ashif achieves a Cost of 4.7% while with the MR value of 12.4%. Fine-tuning the models improves performance on the test documents; however, the patterns, which are more complex, in the unseen test documents, are still not fully captured by $\mathcal{K}$ashif.
To summarize, $\mathcal{K}$ashif consistently outperforms ST29 across all documents, with a notable difference in the number of partial matches. Results thus suggest that fine-tuning pre-trained models on a dedicated dataset is indeed necessary for the model to learn about the LRT task. However, while better than the pre-trained model, $\mathcal{K}$ashif shows the following limitations: 1) it does not provide a rationale behind selecting a trace link, except for the fact that semantic similarity exceeds a pre-defined threshold. This is expected to impede its practical use. 2) The average success rate achieved by $\mathcal{K}$ashif is about 44%, which is not particularly effective.
The answer to RQ3 is that $\mathcal{K}$ashif outperforms ST29, demonstrating that fine-tuning helps the model learn about the LRT task. However, the performance of $\mathcal{K}$ashif still shows significant room for improvement in unseen domains.
4.8 Effectiveness of Large Language Model (RQ4)
The baselines’ performance on the LRT task is extremely poor, highlighting the need for improvement (RQ2). When we attempted a more refined approach using the ST models ($\mathcal{K}$ashif), it performed better than the baselines but did not achieve satisfactory results on an unseen dataset (RQ3). This indicates that the ST models can partially address some of the issues inherent in the baseline approaches. However, the ST models lack the robustness needed to generalize effectively across unseen data, as discussed in RQ3. Given their promising results across many tasks AroraHH24 ; abs_2310_18648 , RQ4 aims to assess whether prompting LLMs with an engineered prompt offers a meaningful alternative for LRT and to compare its effectiveness against prompts that were not engineered. We posit that LLMs, with their pre-training across different domains, might significantly improve trace link recovery.
Metholodology. As discussed in Section 3.3, we designed a prompt, based on the Rice structure vogelsang2024using . We prompted the GPT-4o model to generate recommendations for trace links between requirements and GDPR provisions. We base our analysis on the four documents discussed in RQ3. We compare the LLM’s recommendations, generated with our prompt for each requirement, against our ground truth. We begin by evaluating Rice_LRT under two temperature values (0 and 0.5) and five different random seed values. Then, we compare its performance to the baseline prompts using their reported best settings.
Table 8: Rice_LRT performance across different seed and temperature values.
[t] seed Temperature=0 Temperature=0.5 TP FP F2 EM PM SR MR C TP FP F2 EM PM SR MR C 0 RD1 46 108 58.4 6 55 83.6 77.5 7.9 43 115 55.7 5 54 80.8 77.0 8.3 RD2 52 134 58.3 2 53 85.9 83.8 11.1 49 126 56.3 4 46 78.1 72.5 10.5 RD3 38 112 59.0 2 54 91.8 88.6 9.4 35 111 55.0 3 50 86.9 80.0 9.2 RD4 76 150 66.7 6 79 92.4 92.1 9.4 74 149 65.3 8 75 90.2 89.3 9.3 Average - - 60.6 - - 88.4 85.8 9.4 - - 58.1 - - 84 79.7 9.3 16 RD1 45 107 59.2 5 56 83.6 75.0 7.9 40 111 52.8 6 51 78.1 72.5 7.9 RD2 52 124 59.6 4 51 85.9 78.4 11.1 47 127 54.1 3 46 76.6 73.5 10.4 RD3 38 109 59.6 3 53 91.8 87.1 9.1 35 115 54.3 0 53 86.9 82.8 9.4 RD4 76 157 65.9 4 81 92.4 90.7 9.6 74 141 66.2 7 76 90.2 89.3 8.9 Average - - 61.1 - - 88.4 82.8 9.4 - - 56.9 - - 82.9 79.5 9.1 42 RD1 41 110 54.1 6 51 78.1 72.5 7.9 46 105 60.7 6 56 84.9 85.0 7.9 RD2 52 131 58.7 3 52 85.9 82.3 10.9 46 124 53.5 3 46 76.6 71.5 10.2 RD3 38 106 60.1 3 53 91.8 88.5 9.0 36 112 56.3 2 52 88.5 85.7 9.3 RD4 75 156 65.2 4 80 65.2 90.7 9.6 75 138 67.3 10 74 91.3 90.7 8.9 Average - - 59.5 - - 80.3 83.6 9.3 - - 59.5 - - 85.3 83.2 9.1 256 RD1 42 105 56.0 5 53 79.5 75.0 7.7 39 112 51.5 5 51 76.7 70.4 7.9 RD2 52 135 58.2 3 51 84.4 81.8 11.2 49 137 54.9 1 50 79.7 74.0 11.1 RD3 38 110 59.4 3 53 91.8 88.6 9.3 37 108 58.4 1 54 90.2 87.1 9.1 RD4 76 156 66.0 4 81 92.4 91.3 9.6 74 162 63.8 7 76 90.2 89.3 9.8 Average - - 59.9 - - 87.0 84.2 9.4 - - 57.1 - - 84.2 80.2 9.5 13374224 RD1 41 107 54.4 5 52 78.1 72.5 7.7 39 106 52.3 5 50 75.3 71.2 7.6 RD2 52 129 59.0 3 52 85.9 83.8 10.8 52 132 58.6 3 52 85.9 82.3 11 RD3 37 112 57.6 2 53 90.2 87.1 9.3 38 103 60.7 3 53 91.8 88.6 8.8 RD4 76 153 66.3 4 81 92.4 92.1 9.5 73 157 63.6 5 77 89.1 88.7 9.6 Average - - 59.3 - - 86.7 83.9 9.3 - - 58.8 - - 85.5 82.7 9.2
Evaluation Metrics. Same as in RQ3.
Results. First, to investigate how Rice_LRT performs under different random seed and temperature settings, we run the prompt with five distinct random seeds and two temperature values (0 and 0.5), the latter of which introduces additional randomness in the response. In other words, raising the temperature increases the model’s output entropy, meaning it becomes more likely to select lower-probability tokens. Table 8 shows the results of the Rice_LRT approach, by prompting GPT-4o, using these settings. We observe that the average performance changes slightly when the temperature is held constant and the seed value is changed. The maximum changes are 2.6 pp in F2 score (the temperature is 0.5, and the seed changes from 16 to 42) and 8.1 pp in SR value (the temperature is 0, and the seed changes from 0 to 42).
In contrast, increasing randomness by setting the temperature equal to 0.5 caused a drop in F2 score for a constant seed value. The highest drop was observed for seed 16, where the F2 score dropped from 61.1% to 56.9%. Moreover, we observe that the model’s behavior is not consistent across test documents. For example, when using seed value 13374224, increasing the temperature from 0 to 0.5 results in a decrease in F2 score for RD1, RD2, and RD4, while it leads to an improvement for RD3. Nonetheless, we observe that Rice_LRT shows little variation across different seeds when the temperature is set to 0, suggesting that random variation in results is primarily due to higher temperature. Thus, we select the performance using temperature 0 and arbitrarily select seed 16 to compare against the baselines.
Table 9 shows the results of the Rice_LRT approach and the baseline prompts. Rice_LRT outperforms all prompts at both the trace and requirement levels. P1 includes a retrieval phase that selects the top-k artifacts based on semantic similarity. In our case, the artifacts are the 26 GDPR provisions. Therefore, we experimented with two k values: 13 and 26, corresponding to half and all the GDPR provisions, respectively. Since using k=13 resulted in a 12.6% loss of provisions during the retrieval step by the semantic similarity technique, we chose to include all provisions, given that the total number of provisions is not excessively large in our case. P1 exhibits a lower number of retrieved links (TP+FP) compared to Rice_LRT and P2. Based on our observations of the rationale generated by the model, this stems primarily from its inability to capture indirect relationships between requirements and their corresponding provisions. For instance, a requirement stating that a user’s profile must include at least one charging point may seem purely functional but implicitly involves handling personal data, such as user identity or location. Consequently, it should be linked to the GDPR provision on obtaining user consent for data processing. The prompt’s inability to capture this implicit connection results in lower retrieval performance and the omission of true trace links. Moreover, P1’s F2 score is only 0.8 pp higher than $\mathcal{K}$ashif (13.5% in Table 7), which suggests that this prompt does not fully leverage GPT-4o’s capabilities, while $\mathcal{K}$ashif already captures much of the underlying knowledge by being tuned over another dataset, allowing it to remain highly competitive. P2 exhibits better performance, achieving a higher retrieval compared to P1. We observed that the rationales generated by P2 provide more details in explaining how to link specific requirements to the relevant provisions. While P2’s high retrieval rate increased the number of TPs, it concurrently increased the FPs. This ultimately led to a higher F2 score than P1. P3_1 is similar to P2 in terms of a low retrieval rate, while P3_2 resulted in producing the highest FPs compared to all other prompt templates. Moreover, $\mathcal{K}$ashif has a higher recall compared to P3_1 and P3_2 has a lower F2 score compared to $\mathcal{K}$ashif. On the other hand, the inclusion of context and detailed instructions in Rice_LRT yields more effective outputs from GPT-4o, leading to improved performance at both the trace link and requirement levels. It is worth noting that Rice_LRT, under all temperature and random seed configurations reported in Table 8, consistently outperforms the baselines. Nonetheless, an investigation into which specific elements of the prompt design influence the model’s behavior falls outside the scope of the current study.
Table 9: Accuracy of Rice_LRT approach and baselines on the test datasets (RQ4).
[t] Trace Link Level Requirement Level N T TP FP R P F2 EM PM SR MR C Rice_LRT RD1 73 57 45 107 78.9 29.6 59.2 5 56 83.6 79.2 8.4 RD2 64 65 52 124 80.0 29.5 59.6 4 51 85.9 82.4 10.6 RD3 61 43 38 109 88.4 25.9 59.6 3 53 91.8 88.6 9.2 RD4 92 86 76 157 88.4 32.6 65.9 4 81 92.4 92.3 9.7 Average - - - - 83.9 29.4 61.1 - - 88.4 84.1 9.5 P1 hey2025requirements RD1 73 57 16 25 28.1 39.0 29.7 31 5 49.3 18.7 2.1 RD2 64 65 15 46 23.3 24.6 23.4 31 6 57.8 23.0 3.7 RD3 61 43 1 8 2.3 11.1 2.8 25 0 40.9 1.4 0.5 RD4 92 86 1 9 1.2 10.0 1.4 17 0 18.5 0.6 0.04 Average - - - - 13.7 21.2 14.3 - - 41.6 13.4 1.6 P2 ronanki2024requirements RD1 73 57 38 235 67.9 13.9 38.2 3 52 75.3 65.4 14.4 RD2 64 65 45 475 69.2 8.6 28.8 0 49 76.6 71.0 31.2 RD3 61 43 19 291 44.2 6.1 19.7 0 37 60.7 19.5 19.5 RD4 92 86 36 525 41.9 6.4 19.9 1 42 46.7 37.0 23.4 Average - - - - 55.8 8.7 26.6 - - 64.8 48.2 22.1 P3_1 ge2025cross RD1 73 57 18 10 31.6 64.3 35.2 36 4 54.8 28.3 1.5 RD2 64 65 9 30 13.8 32.1 15.1 29 3 50.0 13.2 2.4 RD3 61 43 5 11 11.6 31.2 13.3 26 1 44.3 8.5 1.0 RD4 92 86 1 4 1.2 20.0 1.4 19 1 21.7 0.6 0.2 Average - - - - 14.5 32.3 16.2 - - 42.7 12.6 1.3 P3_2 ge2025cross RD1 73 57 26 445 45.6 5.5 18.6 1 47 65.7 43.3 24.8 RD2 64 65 23 465 35.4 4.7 15.4 0 38 59.4 38.2 29.3 RD3 61 43 7 413 16.3 1.7 5.9 1 29 49.1 14.3 26.4 RD4 92 86 16 532 18.6 2.9 8.9 1 33 37.0 18.7 22.9 Average - - - - 29.0 3.7 12.2 - - 52.8 28.6 25.8
  • T: Predicted trace links, EM: Exact Match, PM: Partial Match, SR: Success Rate, MR: Macro Recall, C: Cost.
To more effectively illustrate how Rice_LRT and the baselines perform on a per-requirement basis, we report their MR and Cost values. Figure 6 depicts the performance of each prompt. The vertical position of each circle indicates the average recall for retrieving trace links for an individual requirement, while the circle’s radius reflects the total number of retrieved links, including both TPs and FPs. As visible, on average, Rice_LRT achieves the highest recall at 84% while retrieving only 9.5% of all provisions.
Figure 6: MR and Cost comparison on the test documents (GDPR). The y-axis represents the average rate of successful trace-link predictions per requirement, while the bubble size indicates the total number of retrieved trace links. The larger the bubble, the more items retrieved from the entire set.
At the requirements level, the number of exact matches is very low for Rice_LRT. We note that Rice_LRT outputs at least one regulatory code for each requirement (based on our prompt of Section 3.3) even when requirements do not have any trace links in the ground truth. This is one explanation for the sharp decrease in exact matches. Despite this, the number of partial matches has increased significantly, thereby improving the overall success rate. While one would ideally like an approach with a high exact-match rate, we note that the results remain beneficial, as we discuss next.
Fig. 7 shows the split of partially matched requirements for the number of FPs. For instance, for RD1, there were 56 partially matched requirements. Of these, 24 (42.9%) had only one FP, 25 (44.6%) had two FPs, and the remaining 7 (12.5%) had three FPs. As seen in the figure, across all four documents, there were very few requirements with four FPs. This indicates that most partially matched requirements had a manageable number of FPs, typically ranging from one to three. This result is significant because it suggests that the model’s outputs are not overwhelming for analysts to process. Fewer FPs per requirement allow analysts to efficiently review and validate the suggested trace links, reducing their cognitive load. Instead of starting from scratch or sifting through a vast set of 26 possible provisions per requirement, analysts can focus on validating and refining a much smaller, pre-filtered set of trace links. This aligns with the principle of assisted decision-making skitka1999does , in which automated tools augment human judgment by narrowing options.
Our results further indicate that the GPT-4o model successfully demonstrated a stronger understanding of the LRT task when provided with an engineered prompt. This indicates that Rice_LRT is effective at identifying the underlying logic and rationale behind provisions, even with a limited number of few-shot examples. Its ability to navigate complex relationships and extract logical links demonstrates its robustness in understanding the nuances of regulatory requirements. However, the cases it misses highlight areas where the connections may require deeper domain-specific knowledge or additional context to resolve ambiguities.
On investigating the FPs for each requirement, we observed that several predicted trace links may be relevant depending on the application context, even though they do not exactly match the ground truth. These false positives include provisions that are not in the ground truth but are relevant to the input requirements, or provisions in the ground truth (but not all of them), which are why they are considered partial matches. This underscores the potential of Rice_LRT to identify trace links that correspond to potential associations between requirements and provisions that may not have been contemplated when building the ground truth. Such cases could still be informative to the analysts. For example, the Rice_LRT output presented in Section 3.3 included three predictions with corresponding rationales. Of these, [SEC] is the ground truth, and [ACC] and [CNF] are categorized as FPs. The rationale for [ACC] highlights that requiring a key file ensures proper authentication, which can be interpreted as supporting the right to access. Similarly, the rationale for [CNF] emphasizes that protecting the database with a key file ensures sensitive data remains confidential. While these codes are not explicitly part of the ground truth for this requirement, they surface related regulatory considerations that may enrich the analyst’s understanding of the requirement and its broader implications under GDPR. Hence, while FPs may not align perfectly with the ground truth, their contextual relevance, as inferred from the generated rationale, can offer valuable insights for the LRT task. This also underscores the inherent subjectivity of the LRT task, especially when dealing with broadly framed regulations like GDPR, which often leave room for interpretation, compared to domain-specific regulations such as HIPAA.
Figure 7: Number of FPs for requirements with partial match produced by Rice_LRT (Keepass: RD1, WASP: RD2, Datahub: RD3, and ScrumAlliance: RD4).
To assess the statistical significance of the differences in proportions of TPs versus FNs (recall) and TPs versus FPs (precision) between $\mathcal{K}$ashif and Rice_LRT, we employ Fisher’s exact test fisher1922interpretation . This test calculates the probability of obtaining the observed data (TPs, FPs, and FNs) under the null hypothesis that proportions are equal across techniques. The p-values for RD1 for the proportions of (TPs, FPs) and (TPs, FNs) are 3e-4 and 1e-5, respectively. For RD2, the corresponding values are 2.5e-2 and 1e-5, for RD3, 4.6e-3 and 1e-5, and for RD4, 0 and 1e-5. The comparison between Rice_LRT and P1, , P3_1, and P3_2 yields values comparable to those reported for $\mathcal{K}$ashif, while the results for P2 are presented as follows: 0 and 5e-1, 1e-4 and 6e-1, 1e-4 and 5e-2, and 1e-4 and 3e-2 for RD1, RD2, RD3, and RD4, respectively. As we can see, across the four test documents, the resulting p-value is very small (below 5e-2). It falls well below the usual significance levels, indicating that the differences in precision and recall among $\mathcal{K}$ashif, P1, P2, P3_1, , P3_2 and Rice_LRT are statistically significant in most cases, except for P2 and the recall scores for RD1 to RD2.
The answer to RQ4 is that our Rice-based approach, which utilizes an engineered prompting on GPT-4o, outperforms the baselines on the LRT task across the four test documents. Further, the engineered context and instructions used in the few-shot setting of Rice_LRT achieved the highest F2 score by maximizing TPs while reducing FPs compared to the baselines. It also significantly outperforms $\mathcal{K}$ashif, although $\mathcal{K}$ashif remains competitive against P1 and P3_2 prompts.
4.9 Discussion
We experimented with a variety of classifiers, including two IR-, one ML-, two LM-, and one LLM-based methods. We compared them against our ST-based classifier, $\mathcal{K}$ashif, on the LRT task using the publicly available HIPAA dataset. The results show that $\mathcal{K}$ashif achieves the highest F2 score among all baselines, including B, TraceBERT, and LLaMA, which have been previously proposed for inner-requirements traceability. Moreover, $\mathcal{K}$ashif achieved a strong overall ranking performance, obtaining an MAP of 81% and the highest discrimination across thresholds, with an AUC of 0.93. $\mathcal{K}$ashif demonstrated a superior ability to address the linguistic discrepancies between requirements and provisions, outperforming existing models in the literature that target traceability across different artifact types. We also demonstrated that Rice_LRT, as an engineered prompt, surpasses baseline prompts introduced in prior work, achieving a recall of 84%, a precision of $\approx$30%, and F2 score of 61% on a newly created, more complex dataset using the GDPR provisions. From a more practical standpoint, we analyzed Rice_LRT ’s performance per requirement, using both Macro Recall (MR) and Cost. Rice_LRT achieves an MR of 84% with a Cost of 9.5%, meaning that, on average, only 9.5% of the total provisions are retrieved, while ensuring that 84% of the true trace links are included in the retrieved list. This represents a 36 pp improvement in MR over the best baseline prompt, while reducing Cost by 13 pp. Moreover, we show that simple Yes/No–style prompts achieve low MR and Cost retrieval on the test documents, indicating that they do not fully leverage GPT-4o’s capabilities. Lastly, our $\mathcal{K}$ashif classifier achieved competitive F2 scores on our newly created GDPR dataset compared to P1 and P3_2 prompts, despite being fine-tuned only on the HIPAA dataset.
5 Threats to Validity
Internal Validity. Bias is a well-known internal validity concern. To mitigate bias, in RQ3 and RQ4, the dataset of over four documents was curated by two annotators with more than a decade of experience in RE. Before the traceability sessions, there was no exposure to the technical details of our approach. The second potential threat to internal validity concerns the few-shot prompting in RQ4. The initial few-shot examples used for GPT-4o’s prompt engineering could introduce confirmation bias, potentially influencing the model’s predictions. To mitigate this, we designed the few-shot examples to reflect realistic usage scenarios where LLM is a recommendation tool guided by a human expert’s rationale for the first few requirements. This approach aligns with practical applications while minimizing the risk of confirmation bias. Additionally, the limited number of examples in the few-shot prompt was deliberately chosen to avoid overfitting. By doing so, we allowed the LLM sufficient flexibility to independently apply reasoning across the remaining requirements, maintaining a balance between guidance and adaptability. This approach ensures the LLM’s outputs remain broadly applicable while minimizing potential validity threats, as seen by the relatively high success rate in RQ4.
External Validity. We evaluated $\mathcal{K}$ashif on two datasets, namely HIPAA and four new documents against GDPR. HIPAA is a pre-existing dataset frequently used in the RE literature. The test documents used in RQ3 and RQ4 (four new documents against GDPR), which we created as part of our work, covers two types of textual requirements, including user stories and shall-type requirements. Such diversity helped increase the generalizability of our results. Experiments on more diverse requirements documents and other regulations are nonetheless required to improve the external validity of our study.
6 Related Work
Requirements traceability (RT) has been extensively studied in RE mucha2024systematic ; li2023applications ; tufail2017systematic ; wang2018requirements ; ramesh1998factors . Existing work applies different technologies, ranging from traditional methods such as Information Retrieval (IR) and statistical models to more advanced approaches like Machine Learning (ML) and Deep Learning (DL). Early works borrowed IR techniques such as Vector Space Models (VSM), Latent Dirichlet Allocation (LDA), to find trace links between software artifacts via text relevancy gao2022using ; guo2017tackling ; kuang2017analyzing ; nishikawa2015recovering ; panichella2013and ; Mahmoud2013 ; Wang20193 ; sun2017frlink ; capobianco2013improving ; chhabra2017filtering ; bavota2014enhancing ; shao2013improved ; Wang20191 ; Li2020 ; moran2020improving . More advanced techniques have been introduced using ML Rasiman2022 ; Bella2019 ; mills2019tracing ; Bella2018 ; Mills2018 ; Rath2018 ; Zhao2018 ; Mills2017 ; Mills20172 ; Falessi2017 ; Hayes2015 ; le2015rclinker ; li2015recovering and DL kenton2019bert ; Zhang2021 ; Wang20192 ; Guo:17 ; Chen2019 ; Alhoshan2019 ; Alhoshan2018 ; Li2018 ; zhao2020extended ; Sultanov2013 , employing various algorithms — from classifiers like SVM, random forest, and decision trees to more sophisticated language models like BERT devlin-etal-2019-bert to find trace links. In recent years, with the emergence of LLMs, researchers have leveraged pre-trained knowledge through prompt engineering techniques to identify trace links between software artifacts hassine2024llm ; rodriguez2023prompts ; vogelsang2025impact . Hassine hassine2024llm proposed an LLM-based technique that uses zero-shot learning on GPT3.5 to find trace links between requirements and goals in Goal-oriented Language (GRL) models. Moreover, Rodriguez et al. rodriguez2023prompts proposed an approach that integrates zero-shot prompting with reasoning to enhance results in the Traceability Link Recovery (TLR) problem on diverse software artifacts. They have shown that a prompt that performs well with one model or dataset may not yield optimal results with another, underscoring the need to tailor prompts to the specific context. Recent studies have explored the use of prompting techniques for traceability within requirements fuchss2025lissa ; ronanki2024requirements ; fuchss2025beyond ; hey2025requirements . Hey et al. hey2025requirements used Retrieval-Augmented Generation (RAG) with LLMs for inter-requirements traceability, where the model first retrieves relevant candidate requirements. Then, it identifies trace links among the retrieved candidates using the KISS prompt, which consists of a yes/no question prompting the model to determine whether a trace link exists between a given pair of artifacts. Fuchß fuchss2025beyond et al. proposed an ensemble of LLM-based prompts for candidate filtering, where multiple LLM models are iteratively used to re-rank or filter the initial set of retrieved candidates. The prompt template used in this work follows the same design as that proposed by Hey et al. hey2025requirements . Additionally, Ronanki et al. ronanki2024requirements developed and experimented with five distinct interactive prompting patterns, each tuned for a specific dataset.
In addition to the algorithms being used, the types of artifacts with which these algorithms are intended to work also play a significant role. Existing studies primarily focus on identifying trace links between requirements and code north2024code ; peng2023enhancing ; cleland:2010 ; panichella2013and ; kuang2017analyzing ; gao2022using ; vogelsang2025impact . Only a few studies have focused on establishing traceability across different software artifacts nishikawa2015recovering . Existing approaches for RT are not directly applicable in our context due to the significant discrepancy between the legal language used in regulations and the technical language used in software requirements and related artifacts.
Legal requirements traceability has only been investigated to a limited extent in the literature. Cleland-Huang et al. cleland:2010 propose a probabilistic approach that identifies trace links between requirements and the HIPAA regulation by computing probability values based on detecting indicator terms for regulations that correspond to requirements. The authors further propose extending the indicator terms with more domain-specific terms retrieved from the web. In a follow-up work, Gibiec et al. Gibiec:2010 further investigate mining the web. Guo et al. Guo:17 extend the previous two papers to improve the terminology gap problem, i.e., the mismatch between terms in requirements and regulations. The authors investigate different methods based on classification, ontologies, and web-mining and evaluate their approaches on HIPAA.
While previous research has made significant strides in requirements traceability using traditional IR methods and ML/DL techniques, these approaches exhibit notable limitations in addressing the complexities of the LRT task. Most notably, existing methods struggle with the terminology gap between regulations and technical requirements, do not generalize well across regulations, and lack adaptability to multi-domain applications. Additionally, studies that have proposed prompt-based approaches have not fully leveraged the capabilities of prompt engineering, providing only detailed instructions and limited context. Their approaches primarily rely on pre- and post-processing strategies hey2025requirements ; fuchss2025beyond , which in some cases query several large language models. In comparison to the above works, we empirically evaluate two automated LRT approaches: (1) a classifier-based solution leveraging sentence transformers and (2) a generative LLM-based solution guided by structured prompt engineering. By exploring these methods across two distinct regulations, HIPAA and GDPR, we advance the understanding of how modern NLP techniques can be adapted to meet the challenges of LRT. We also shed light on the possibilities, or lack thereof, of transfer learning across regulations. To the best of our knowledge, we are also among the first to identify the strengths and limitations of LLMs in this context. Further and larger studies with human experts are required to establish the benefits of LLMs for LRT.
7 Conclusion
This study presents a comparative evaluation of two approaches for Legal Requirements Traceability (LRT): a classifier-based method, $\mathcal{K}$ashif, leveraging sentence transformers, and a generative LLM-based method, Rice_LRT, designed using a structured prompt engineering framework. Moreover, we compare $\mathcal{K}$ashif against seven distinct baselines—ranging from traditional IR techniques to recent deep learning and transformer-based models (LSI, LDA, GloVe, TraceBERT, RoBERTa, and LLaMa2)—to comprehensively demonstrate its capabilities. Our results demonstrate that $\mathcal{K}$ashif provides significant improvements over the baselines in terms of F2 score, achieving a F2 score of $\approx$63% on HIPAA data (23 pp more than the best baseline). On the other hand, $\mathcal{K}$ashif’s performance deteriorates on more complex datasets such as GDPR, yielding only 15% recall. However, this also shows that $\mathcal{K}$ashif, despite being tuned on a similar but less complex dataset, achieves performance comparable to a simple prompt (answering yes/no) on GPT-4o, which does not fully exploit its capabilities.
Conversely, the Rice_LRT approach, built on generative LLMs, outperformed $\mathcal{K}$ashif on GDPR data, achieving an F2 score of 61% and reducing manual effort for traceability by enabling analysts to vet only a fraction of trace links. Moreover, Rice_LRT performed better than a simpler prompt, increasing TPs while reducing FPs. These findings suggest that generative LLMs and carefully designed prompts offer a promising approach to automating LRT tasks in complex legal domains. However, the approach has its challenges, including the occurrence of false positives that require further investigation. In addition to evaluating current state-of-the-art methods, this work highlights critical challenges, including terminology gaps between requirements and provisions By addressing these challenges, our study underscores the importance of tailoring solutions to the nuances of legal and regulatory contexts.
In the future, we plan to conduct a human-in-the-loop study with a domain expert to investigate the applicability of LLMs in the LRT context. We further plan to enhance LLM performance by incorporating domain-specific knowledge to better address terminology and contextual gaps between regulatory texts and technical requirements, particularly in the GDPR context.
Acknowledgements
The experiments conducted in this work were enabled by support provided by the Digital Research Alliance of Canada.
Declarations
Authors Contributions Romina Etezadi: Conceptualization, Investigation, Formal analysis, Software, Writing - original draft. Sallam Abualhaija: Conceptualization, Investigation, Data Curation, Formal analysis, Writing - original draft & review & editing. Chetan Arora: Conceptualization, Investigation, Data Curation, Formal analysis, Writing - review & editing. Lionel Briand: Conceptualization, Investigation, Formal analysis, Supervision, Writing - review & editing, Funding acquisition.
Funding This work is supported by the H2020 COSMOS European project, grant agreement No. 957254, NSERC of Canada under the Discovery and CRC programs, the Research Ireland grant 13/RC/2094-2, and the Luxembourg National Research Fund under grant number C23/IS/17958091/PLAITO. It is part of a collaborative research program between the University of Ottawa’s Nanda laboratory and the SnT centre at the University of Luxembourg.
Conflict of Interest The authors declared that they have no conflict of interest.
Ethical Approval Not applicable.
Informed Consent Not applicable.
Clinical Trial Number Not applicable.

### v6

@@
 4 Lero Research Ireland centre for Software Research and University of Limerick, Ireland
-Classification or Prompting: A Case Study on Legal Requirements Traceability
+Classifier or Prompt: A Case Study on Legal Requirements Traceability
 Romina Etezadi1    Sallam Abualhaija2    Chetan Arora3    Lionel Briand1,4
@@
 This evolution, driven by recent breakthroughs in artificial intelligence (AI), has led to increasing complexity in software systems ahmad2023requirements ; Feldt:18 .
-As technology progresses, regulations are adapting in parallel to ensure that software systems are developed in line with ethical and legal standards.
-For example, the general data protection regulation (GDPR) GDPR is enforced since 2018 to address concerns about privacy and data protection.
-Despite being introduced by the European Union (EU), the GDPR has a global effect, impacting organizations (and software) outside the EU as long as they handle personal data of EU residents.
+As technology advances, regulations are evolving in parallel to ensure that software systems are developed in accordance with ethical and legal standards.
+For example, the General Data Protection Regulation (GDPR) GDPR has been in effect since 2018 to address concerns regarding privacy and data protection.
+Although introduced by the European Union (EU), the GDPR has a global effect, affecting organizations (and software) outside the EU that process personal data of EU residents.
 Requirements Engineering (RE) plays a pivotal role in this landscape.
@@
 One scenario is to explicitly identify legal requirements early during the requirements elicitation phase, answering the question: “What legal obligations need to be satisfied by the system for it to be compliant?”.
-The elicited legal requirements can then be integrated into the software development process, while maintaining trace links to the source legal provisions.
-As an alternative scenario, requirements engineers may need to verify the compliance of existing software systems against legal provisions in a post-deployment stage, as new regulations have become applicable.
+The elicited legal requirements can then be integrated into the software development process while maintaining traceability to the source legal provisions.
+As an alternative scenario, requirements engineers may need to verify compliance of existing software systems with legal provisions during the post-deployment stage, as new regulations have become applicable.
 In this case, they must answer the question “Does the system satisfy the regulation?”.
-To do so, engineers must analyze the regulation, identify applicable legal provisions, and then trace software requirements to these statements.
-Both alternatives rely on requirements traceability analysis, an essential RE activity concerned with the identification and maintenance of trace links between requirements and other artifacts within the software development lifecycle Meyer:22 .
+To do so, engineers must analyze the regulation, identify the applicable legal provisions, and trace software requirements to those provisions.
+Both alternatives rely on requirements traceability analysis, an essential RE activity concerned with identifying and maintaining trace links between requirements and other artifacts throughout the software development lifecycle Meyer:22 .
 legal requirements traceability (LRT) is a special case where requirements are traced to provisions in a regulation and is the focus of this paper.
 To illustrate this concept, consider the following example.
-Imagine a fictional mobility app named WeMobilize, which helps users book and share cab rides.
-Originally a non-EU startup, WeMobilize is expanding to the EU and hence must comply with the GDPR.
-This example is particularly relevant as many businesses are globalizing and must adapt to data protection laws in different jurisdictions.
+Consider a fictional mobility app named WeMobilize, which helps users book and share taxi rides.
+Originally a non-EU startup, WeMobilize is expanding into the EU and must therefore comply with the GDPR.
+This example is particularly relevant, as many businesses are globalizing and must adapt to data protection laws across jurisdictions.
 Fig.
-1 shows how WeMobilize’s requirements (labeled REQ1 – REQ5) can be traced to data protection policies in the GDPR GDPR .
-We identify trace links to provisions in GDPR for REQ1 and REQ3 – REQ5, visualized as dashed lines in black.
-REQ2 has no trace link to GDPR in our example since it does not involve processing users’ personal data.
-REQ1 involves collecting user’s personal information and must therefore be traced to two provisions, namely REG_DIR (related to the direct collection of personal information) and REG_CON (related to the explicit soliciting of users’ consent).
-Currently, consent is not part of REQ1, which prevents identifying a trace link with REG_CON—a missing trace link is visualized with a red dashed line in the figure.
-Failing to identify this trace link entails a possible breach of GDPR.
-Therefore, deploying WeMobilize as-is, without accounting for provisions in GDPR, can lead to potential reputational and financial losses caused by violating the GDPR.
-LRT can help identify potential non-compliance issues at early stages but requires not only legal expertise but also substantial manual effort.
+1 illustrates how WeMobilize’s requirements (labeled REQ1–REQ5) can be traced to data protection policies under the GDPR GDPR .
+We identify trace links to provisions of the GDPR for REQ1 and REQ3—REQ5, visualized as dashed black lines.
+REQ2 has no trace link to GDPR in our example, as it does not involve processing users’ personal data.
+REQ1 involves collecting users’ personal information and must therefore be traced to two provisions: REG_DIR (for the direct collection of personal information) and REG_CON (for the explicit solicitation of users’ consent).
+Currently, consent is not part of REQ1, which prevents the identification of a trace link with REG_CON—a missing trace link is visualized as a red dashed line in the figure.
+Failing to identify this trace link may constitute a breach of the GDPR.
+Therefore, deploying WeMobilize as-is, without accounting for GDPR provisions, can lead to reputational and financial losses arising from GDPR violations.
+LRT can help identify potential non-compliance issues at an early stage, but it requires not only legal expertise but also substantial manual effort.
 Developing automated support is therefore beneficial to assist engineers and analysts in identifying applicable trace links.
@@
 Guo:17 focused on bridging the terminology gap between provisions and software requirements.
-They examined three methods, including the one by Cleland-Huang et al.
-mentioned above, and two others based on web-mining and ontologies.
+They examined three methods, including the method by Cleland-Huang et al.
+mentioned above, as well as two others based on web mining and ontologies.
 The proposed methods aim to expand the terminology of the provisions by adding additional terms to better identify trace links.
@@
 Third, prompt engineering has not been given sufficient attention; existing work often relies on overly simple prompts ronanki2024requirements ; hey2025requirements ; ge2025cross or on querying multiple LLMs fuchss2025beyond , which may not always be feasible in real-world deployments due to computational and cost constraints.
-Fourth, the evaluation is based on a single benchmark that may not fully reflect the complexity of the legal domain in practice.
+Fourth, the evaluation relies on a single benchmark that may not fully capture the complexity of the legal domain in practice.
 To address these limitations, we propose two novel approaches in this paper, based on recent NLP technologies that utilize the Transformer architecture Vaswani:17 and LLMs.
-Similar to existing work, both approaches aim to predict trace links, and we assess their performance on a realistic scenario beyond the commonly used benchmark dataset.
+Similar to prior work, both approaches aim to predict trace links, and we evaluate their performance in a realistic scenario beyond the commonly used benchmark dataset.
 Contributions. The paper makes the following contributions:
@@
 More details on this evaluation can be found in Section 4.6.
-(3) To further confirm its performance, we test $\mathcal{K}$ashif on new unseen requirements documents covering diverse domains and requirements types.
-These requirements are traced to the GDPR, a more complex regulation with 26 provisions pertaining to personal data protection that must be adhered to in software requirements.
+(3) We created a new dataset based on GDPR regulations.
+It includes 310 requirements covering diverse domains and requirements types.
+These requirements are traced to the GDPR, a complex regulation comprising 26 provisions on personal data protection, to which software requirements must comply.
+(4) To further confirm $\mathcal{K}$ashif’s performance, we test it on our newly created unseen requirements documents.
 On this dataset, the base ST model (without additional fine-tuning) yields an average recall of 15%.
@@
 For example, an LM would predict “Mat” as the most likely next word in the input sequence, “The cat sits on the [WORD]”.
-LMs are trained on large corpora of texts to estimate these probability distributions accurately.
+LMs are trained on large corpora of text to accurately estimate these probability distributions.
 State-of-the-art LMs are based on transformer architecture, which leverages self-attention mechanisms to weigh the significance of different parts of an input text relative to a given position Vaswani:17 .
-The attention mechanism determines which words in a sentence are more important based on the context and gives them more “attention”.
+The attention mechanism determines which words in a sentence are more important in the context and assigns them greater attention.
 For instance, in the sentence “Mary, who used to live in Paris, loves wine.”, the attention is on Mary and wine.
-Building on transformer architectures, the Sentence Transformers framework (ST) Reimers:19 offers a set of pre-trained models designed to encode longer text sequences, such as sentences or paragraphs, into dense vector representations within a high-dimensional space.
+Building on transformer architectures, the Sentence Transformers framework (ST) Reimers:19 provides a set of pre-trained models that encode longer text sequences, such as sentences or paragraphs, into dense vector representations in a high-dimensional space.
 They produce contextual embeddings that capture the overall semantic essence of an entire input sequence.
-More recently, generative LLMs have emerged as transformer-based language models that are scaled up significantly in model size and the amount of training data.
+More recently, generative LLMs have emerged as transformer-based language models that are significantly larger and trained on much more data.
 Examples on LLMs include OpenAI’s GPT (Generative Pre-trained Transformer) Radford:18 and LLaMa touvron2023llama ; touvron2023llama2 .
@@
 In this paradigm, the ML algorithm is provided with labeled training data where each data point consists of an input vector (features) and the corresponding output label (or value).
-The algorithm learns patterns within the input features to make predictions based on this training.
+The algorithm learns patterns in the input features to make predictions from this training data.
 When trained on a sufficiently large dataset, the algorithm refines its predictions to classify the provided labels more accurately.
-The example ML classification algorithms include random forest, decision tree, support vector machine, and feed-forward neural networks Goodfellow:16 .
+Examples of ML classification algorithms include random forests, decision trees, support vector machines, and feedforward neural networks Goodfellow:16 .
 3 Approaches
@@
 Step 4 preprocesses the input requirements document (RD).
-Step 5 applies the LRT model to compute the semantic similarities between each requirement in the RD and each provision.
+Step 5 applies the LRT model to compute the semantic similarity between each requirement in the RD and each provision.
 Step 6 predicts trace links.
@@
 Step 4: Preprocessing
-In step 4, we preprocess the input requirements using a simple NLP pipeline composed of two modules, namely Tokenization and sentence splitting.
+In step 4, we preprocess the input requirements using a simple NLP pipeline composed of two modules: Tokenization and Sentence Splitting.
 The goal is to decompose the text into separate sentences.
@@
 Below, we discuss alternative methods for setting $\theta$.
-(a) Constant Threshold: To predict a trace link, we utilize a pre-defined constant threshold, $\theta=0.5$.
+(a) Constant Threshold: To predict a trace link, we use a predefined threshold, $\theta=0.5$.
 Specifically, a trace link is predicted if the similarity score exceeds $0.5$.
@@
 Moreover, this threshold was chosen based on the fine-tuning approach used in $\mathcal{K}$ashif.
-During training, positive pairs are labeled with a similarity score of 1, while negative pairs are labeled with 0.
+During training, positive pairs are assigned a similarity score of 1, whereas negative pairs are assigned 0.
 The model is thus optimized to produce similarity scores closer to 1 for true links and closer to 0 for unrelated pairs.
-As a result, a threshold of 0.5 naturally serves as a midpoint between these two extremes, making it a natural choice for distinguishing between positive and negative links.
-(b) Dynamic Threshold: Another practical method to adjust $\theta$ involves curating a set of negative training examples, i.e., requirements that do not have trace links.
+As a result, a threshold of 0.5 serves as the midpoint between these two extremes, making it a natural choice for distinguishing positive from negative links.
+(b) Dynamic Threshold: Another practical method to adjust $\theta$ involves curating a set of negative training examples, i.e., requirements that lack trace links.
 These requirements can be sourced from publicly available datasets or from different projects.
@@
 (c) Maximum Delta Cutoff: In this method, we apply the following procedure.
-First, for each $r_{i}$, we sort the similarity values computed across the different provision $c_{j}\in\mathcal{C}$.
+First, for each $r_{i}$, we sort the similarity values computed across the different provisions $c_{j}\in\mathcal{C}$.
 Then, we compute delta values ($\Delta$) corresponding to the differences between each pair of consecutive similarity values and identify the largest $\Delta$ (i.e., the biggest gap in the computed similarities).
@@
 The optimal threshold with the highest F2 score identified on the training set is then applied to the test set.
-The methods described above result in four variants of $\mathcal{K}$ashif, each determined by how $\theta$ is set.
+The methods described above yield four variants of $\mathcal{K}$ashif, each determined by the value of $\theta$.
 These variants are referred to as $\mathcal{K}$ashif constant, $\mathcal{K}$ashif dynamic, $\mathcal{K}$ashif Δ, and $\mathcal{K}$ashif tuned.
@@
 3.3 Rice_LRT
-Our second proposed approach, Rice_LRT, is composed of two steps as illustrated in Fig.
+Our second proposed approach, Rice_LRT, comprises two steps, as illustrated in Fig.
 3.
-The first step involves designing a prompt that is effective for addressing LRT.
+The first step is to design a prompt effective for addressing LRT.
 The second step then applies the prompt to instruct an LLM to predict trace links.
@@
 Step 1: Prompt Design
-In this step, we designed the prompt following recent best practices reported in the RE literature vogelsang2024using ; vogelsang2024specifications .
+In this step, we designed the prompt in accordance with recent best practices reported in the RE literature vogelsang2024using ; vogelsang2024specifications .
 Fig.
 4 presents our final prompt, obtained through iterative refinements.
-The prompt was used on each requirement in the input RD, with the requirement located at the end of the prompt.
-To design the prompt, we followed the Rice (Role, Instruction, Context, Constraints, Examples) framework, with some small adaptation to the LRT task, as we discuss below.
+The prompt was applied to each requirement in the input RD, with the requirement listed at the end of the prompt.
+To design the prompt, we followed the Rice (Role, Instruction, Context, Constraints, Examples) framework, with some small adaptations to the LRT task, as we discuss below.
 The prompt is structured in the following five elements:
@@
 We omitted the explicit mention of the role to obtain a more general applicability of the prompt.
-For LRT, multiple analysts with different backgrounds are likely involved, e.g., a legal analyst in addition to the requirements analyst.
+For LRT, multiple analysts with diverse backgrounds are likely involved, for example, a legal analyst in addition to a requirements analyst.
 The context corresponds to the text shaded in cyan in Fig.
@@
   • Examples: This element provides a few examples selected from our ground truth.
-The examples should cover different trace links.
-Each example is composed of a requirement and the set of trace links alongside the rationale behind each trace link.
+The examples should cover trace links to different provisions.
+Each example comprises a requirement and the set of trace links, along with the rationale for each trace link.
 We note that the LRT task is complex, as we demonstrate throughout the paper.
@@
 4.
+The few-shot examples are not present in the test data.
   • Instruction: This element provides explicit instructions on how to perform the LRT task.
@@
 4.
-Compared to the original Rice framework, this element contains both the Instruction element combined with the Constraint.
-The reason for this is that both elements are intertwined in our context.
+Compared with the original Rice framework, this element combines the Instruction and Constraint elements.
+This is because the two elements are intertwined in our context.
 The prompt must therefore account for task-specific considerations, explained below.
@@
       The prompt should encourage the LLM to equally consider other provisions, since only a subset of the provisions are explicitly explained via the examples and rationales in the Examples element.
-Ideally, the prompt should present an example on each provision.
-However, this is infeasible since only relevant provisions should be traced to software requirements in a given project.
-For instance, if the legal basis for collecting personal data is the contract, then unlike explicit consent, only certain data subject rights are applicable according to GDPR and must be appropriately implemented in the software.
+Ideally, the prompt should present an example of each provision.
+However, this is infeasible, as only the relevant provisions should be traced to the software requirements for a given project.
+For instance, if the legal basis for collecting personal data is the contract, then, unlike explicit consent, only certain data subject rights apply under the GDPR and must be appropriately implemented in the software.
     • $\bullet$
       The prompt should account for indirect trace links.
-As stated above, the LRT is challenging primarily due to the terminology gap between requirements and provisions.
-We therefore encourage the LLM to use its reasoning capability to identify indirect links, generalizing beyond the provided examples in the prompt.
+As noted above, the LRT is challenging primarily due to the terminology gap between the requirements and the provisions.
+We therefore encourage the LLM to use its reasoning capabilities to identify indirect links and to generalize beyond the examples provided in the prompt.
     • $\bullet$
@@
   • Output Indicator: This element clearly describes the output format, corresponding to the text shaded in violet in Fig. 4.
-We note that prompt templates such as Rice serve as illustrative examples or starting points rather than scientifically validated or universally optimal configurations vogelsang2024specifications ; huang2025prompt .
-Prompt templates are typically task-specific and heuristic in nature.
-Based on our preliminary analysis of Rice, providing the example before instructions and contextualizing the instructions turned out to work better rather than providing the example at the end.
 [Context]I am currently working on a task focused on establishing traceability between software requirements and regulatory codes [^3] Note that we use regulatory codes to mean provisions, since the former was used in the literature cleland:2010 .
@@
 Figure 4: Final Rice_LRT prompt for addressing LRT.
+We note that prompt templates such as Rice serve as illustrative examples or starting points rather than scientifically validated or universally optimal configurations vogelsang2024specifications ; huang2025prompt .
+Therefore, during the refinement process, we randomly selected a separate subset from the few-shot example pool (two out of five examples per document; for a total of eight examples) and used them as inputs for refinement, without including them inside the prompt.
+In each iteration, a requirement text from this subset was used to obtain the model’s response, and the generated rationale was then analyzed to refine the instructions by adjusting constraints, ensuring that the resulting rationales were non-hallucinatory, less ambiguous, and judged as logical by manual evaluation.
+Note that, after refinement, these examples were returned to the few-shot sets and used as few-shot examples during the test stage.
+Moreover, the analysis indicates that presenting the example before the instructions and contextualizing the instructions were more effective than presenting the example at the end.
 Step 2: LLM Querying
@@
 A prerequisite for using our prompt involves creating a few examples that will demonstrate the LRT task to the LLM.
-To effectively trigger the reasoning of the LLM, we built five examples by exposing both the labels (i.e., trace links) and the rationale behind selecting these labels.
+To effectively trigger the reasoning of the LLM, we built five examples by exposing both the labels (i.e., trace links) and the rationale for selecting them.
 The examples are then integrated into the above LRT prompt template.
@@
 How accurate is $\mathcal{K}$ashif compared to an existing baseline on a standard dataset from the literature?
-RQ2 aims to assess the value of utilizing ST as enabling technology for addressing the LRT problem compared to a baseline from the existing literature, which we re-implement in this work.
+RQ2 aims to assess the value of using ST as an enabling technology for addressing the LRT problem, compared with a baseline from the existing literature that we reimplement in this work.
 The baseline is a classifier that leverages the terminology probability distributions to compute the likelihood that a requirement can be traced to a provision, based on the occurrence of some indicator terms within that provision.
@@
 In RQ3, we test $\mathcal{K}$ashif on four different documents, two shall-requirements and two user stories, covering various domains.
-These documents are traced to the GDPR privacy requirements.
-The goal of RQ3 is to investigate the performance of $\mathcal{K}$ashif on a more realistic dataset that captures the complexity of the legal domain.
+These documents are derived from the GDPR’s privacy requirements.
+The goal of RQ3 is to evaluate kashif’s performance on a more realistic dataset that captures the complexity of the legal domain.
 RQ4.
@@
 The provisions are the following: access control (AC), audit controls (AUD), person or entity authentication (PA), transmission security (TS), unique user identification (UUI), emergency access procedure (EAP), automatic logoff (AL), encryption and decryption (SED), encryption (TED), and integrity controls (IC).
-HIPAA consists of 10 requirements documents, all shall-requirements, from the healthcare domain.
+HIPAA comprises 10 requirements documents, all of which are shall requirements, from the healthcare domain.
 In total, the dataset contains 1,891 requirements, of which 243 have trace links.
@@
 Total Links 64 73 51 93 281
-To address RQ3, we create and curate four documents covering different requirements types and domains.
-These documents represent a snapshot of a practical scenario that exemplifies the potential complexity of LRT in practice.
+To address RQ3 and RQ4, we create and curate four documents covering different requirements types and domains.
+These documents provide a snapshot of a practical scenario that exemplifies the potential complexity of LRT.
+Moreover, we limit testing of the prompts and LLMs to unseen datasets to minimize the risk of data leakage.
 For each document, we manually identify trace links between software requirements and a list of 26 provisions derived from GDPR and pertinent to software.
@@
 The test documents originate from four independent sources, each from a different domain, as shown in Table 2.
-It is worth mentioning that the test documents are newly created resources that have not yet been publicly released.
+It is worth mentioning that the test documents, unlike HIPAA dataset, are newly created resources that have not yet been publicly released.
 Although the underlying content may not rely entirely on new concepts, the traceability links between the software projects and the GDPR are new and have not been shared before.
@@
 To evaluate our proposed solutions, we compare them against carefully selected baselines.
-we present the baselines for both $\mathcal{K}$ashif and Rice_LRT, covering their classifier-based and prompt-based variants.
+We present the baselines for both $\mathcal{K}$ashif and Rice_LRT, including their classifier- and prompt-based variants.
 4.3.1 Classifier Baselines
@@
 LSI extends the Vector Space Model by applying Singular Value Decomposition (SVD) baeza1999modern to the term-document matrix, thereby generating a lower-dimensional latent semantic space.
-The dimensionality of this subspace is controlled by a manually tuned parameter, commonly referred to as the k value.
+The dimensionality of this subspace is controlled by a manually tuned parameter, commonly referred to as k.
 By capturing the most significant underlying patterns in term usage, LSI implicitly accounts for frequently co-occurring terms, thereby mitigating the vocabulary mismatch problem commonly encountered in information retrieval-based traceability recovery.
@@
 We re-implement as part of this work the baseline B from the literature cleland:2010 ; Guo:17 .
-B is a probabilistic approach based on occurrences of words in requirements texts and how likely these words are associated with specific provisions.
+B is a probabilistic approach based on the occurrences of words in requirements texts and on how likely these words are to be associated with specific provisions.
 Specifically, B predicts whether a requirement is traced to a provision by identifying keywords (also known as indicator terms) that are present in the requirement.
 Given an input requirement for which the trace link should be predicted, B requires a training set from which likelihood estimates of indicator terms are computed for the input requirement, representing how likely it is to be relevant to a specific regulation.
-The training set comprises provisions, software requirements, and the trace links between them.
+The training set comprises provisions and software requirements, along with the trace links between them.
 During training, indicator terms are identified and weighted for each provision by parsing the textual requirements traced to these statements.
@@
 We implement a technique based on static word embeddings using GloVe, which generates dense vector representations of words by leveraging global co-occurrence statistics from a large corpus, thereby capturing both semantic and syntactic relationships between words.
-In our work, we use GloVe as a trainable embedding layer and as the input layer in a simple neural network architecture (with a single layer to predict the score).
-In this context, the inputs consist of a requirement and a candidate provision, and the output is a probability score computed using a sigmoid activation function.
+In our work, we use GloVe as both a trainable embedding layer and the input layer in a simple neural network architecture (with a single-layer architecture to predict the score).
+In this context, the inputs comprise a requirement and a candidate provision, and the output is a probability score computed via a sigmoid activation function.
 The model is subsequently fine-tuned on the training dataset, allowing it to adapt the pre-trained word embeddings to the specific characteristics of the traceability task.
@@
 we utilize the publicly available implementation of TraceBERT, a transformer-based model that, like our approach, is built on a transformer architecture.
-In this paper, the authors propose three distinct BERT-based architectures (TWIN, Single, and Siamese) for fine-tuning the pre-trained BERT model specifically for the traceability link prediction task.
-Each architecture represents a different way to structure the input and optimize the model for semantic similarity between software artifacts.
+In this paper, the authors propose three distinct BERT-based architectures (TWIN, Single, and Siamese) for fine-tuning the pre-trained BERT model for the traceability link-prediction task.
+Each architecture represents a distinct approach to structuring the input and optimizing the model for semantic similarity between software artifacts.
 Based on reported evaluation results, the Single-TraceBERT architecture outperforms the other variations, and we adopt it as one of the baseline models in this work.
@@
 In this work, they use data augmentation to fine-tune LLaMa2 at different scales (1.1B, 7B, and 13B) and demonstrate that LLaMa2-7B achieves the best performance.
-They demonstrate that the fine-tuned LLaMa outperforms both IR-based and ML-based approaches on six datasets, using a 90/10 train–test split.
+They demonstrate that the fine-tuned LLaMa outperforms both IR- and ML-based approaches across six datasets using a 90/10 train–test split.
 In our study, we evaluate how well this approach performs on the requirement–provision task under a more complex evaluation setting.
@@
 For Rice_LRT, we incorporate four baselines, P1, P2, P3_1, and P3_2, from the literature hey2025requirements ; ronanki2024requirements , which focus on identifying links between high-level and low-level requirements.
-This comparison enables us to assess the effectiveness of our context-rich prompt, Rice_LRT, relative to recent prompt-based techniques reported in the literature for traceability.
+This comparison enables us to assess the effectiveness of our context-rich prompt, Rice_LRT, relative to recent traceability prompt-based techniques reported in the literature.
 P1.
@@
 Their study evaluated four zero-shot prompt variations, ranging from a simple prompt to a more detailed CoT formulation.
-In order to have a variety of prompts in our baselines, we employ the prompt with the best results in their work, we call it P3_1 throughout the paper, along with the CoT-based prompt as it includes more detailed instructions compared to P1 and P2, we call it P3_2.
+In order to have a variety of prompts in our baselines, we employ the prompt with the best results in their work, which we call P3_1 throughout the paper, along with the CoT-based prompt, as it includes more detailed instructions compared to P1 and P2, which we call P3_2.
 We modify the prompt to better fit the LRT task.
 The prompts we used are as follows:
-P3-1: Consider the following scenario where the Requirement text represent requirements of the software system and Regulation represent a GDPR regulation.
+P3-1: Consider the following scenario where the Requirement text represents requirements of the software system and Regulation represents a GDPR regulation.
 Let’s think step by step: 1.
@@
 Regulation: '''{content}'''
-P3-2: Consider the following scenario where the Requirement text represent requirements of the software system and Regulation represent a GDPR regulation.
+P3-2: Consider the following scenario where the Requirement text represents requirements of the software system and Regulation represents a GDPR regulation.
 Answer with Yes or No: Does the regulation trace back to the requirement?
@@
 We implement $\mathcal{K}$ashif in Python 3.8.
-For preprocessing the text, we use the NLTK toolkit (v 3.8.1).
+For text preprocessing, we use the NLTK toolkit (v.
+3.8.1).
 We access the ST pre-trained models through the Hugging Face Transformers library (4.44.0).
@@
 LSI and LDA.
-For LSI and LDA, we utilized the scikit-learn library (v1.7.0), setting the number of components to 50 within a search space of 5, 10, 25, and 50 components.
-For LDA the Document–Topic Dirichlet Prior ($\alpha$) and Topic–Word Dirichlet Prior ($\beta$) had the values of 2e-2 within a search space of 1e-2, 2e-2, and 1e-1.
-For both models, we employed TF-IDF as the feature weighting scheme using scikit-learn library (v1.7.0).
+For LSI and LDA, we utilized the scikit-learn library (v1.7.0), setting the number of components $(n_{\text{components}}$) to 50 within a search space of 5, 10, 25, and 50 components.
+For LDA, the Document–Topic Dirichlet Prior ($\alpha$) and Topic–Word Dirichlet Prior ($\beta$) had the values of 2e-2 within a search space of 1e-2, 2e-2, 1e-1, and 2e-1.
+The ranges for $\alpha$ and $\beta$ were chosen to encompass commonly used settings in the literature uto2017diverse ; silva2021topic , with particular emphasis on values proportional to $1/n_{\text{components}}$.
+For both models, we employed TF-IDF as the feature weighting scheme using the scikit-learn library (v1.7.0).
 Our experiments showed that the LSI model achieved optimal performance when stopwords and punctuation were removed.
-In contrast, the LDA model performed best when stopwords and punctuation were removed and lemmatization was applied using NLTK toolkit (v3.8.1).
-We did not apply minimum cutoff thresholds for the vocabulary, as filtering out low-frequency terms could risk omitting important information relevant to some regulatory provisions.
+In contrast, the LDA model performed best when stopwords and punctuation were removed, and lemmatization was applied using the NLTK toolkit (v3.8.1).
+We did not apply minimum cutoff thresholds for vocabulary, as filtering out low-frequency terms could risk omitting information relevant to certain regulatory provisions.
 For the maximum document frequency cutoff, we experimented with values of 0, 1e-1, and 1.
@@
 The best-performing hyperparameters were using epochs=10, output dimension=300, and max_length=50.
-TraceBERT. For TraceBERT, we used the publicly available code from data preparation for tuning the model and retained the fine-tuning settings in its shared implementation.
+TraceBERT. For TraceBERT, we used the publicly available data-preparation code to tune the model and retained the fine-tuning settings from the shared implementation.
 RoBERTa.
 For RoBERTa, the entire fine-tuning setup is implemented using the Hugging Face Transformers library.
-We performed grid search to tune the hyperparameter, considering using epochs=[5, 10, 15], batch size=[8, 16, 32], optimizer=[AdamW, RMSprop], learning rate=[1e-5, 2e-5, 1e-3, 2e-3], and weight decay=[0, 1e-5, 1e-3, 1e-2, 1e-1].
+We performed a grid search to tune the hyperparameters, considering using epochs=[5, 10, 15], batch size=[8, 16, 32], optimizer=[AdamW, RMSprop], learning rate=[1e-5, 2e-5, 1e-3, 2e-3], and weight decay=[0, 1e-5, 1e-3, 1e-2, 1e-1].
 The best-performing hyperparameters were epochs=5, batch size=32, optimizer=AdamW, learning rate=2e-5, and weight decay=1e-2.
@@
 We relied on their released code to generate the training set following the procedure outlined in their paper.
-Since using the same hyperparameters resulted in a single label being produced for all pairs, we performed another round of hyper-parameter tuning considering learning rate=[2e-8, 2e-7, 2e-5, 2e-5, 2e-4, 2e-3, 2e-2] and weight decay=[0, 1e-5, 1e-3, 1e-2, 1e-1], which the best-performing hyperparameters were learning rate=2e-4 and weight decay=1e-5.
+Since using the same hyperparameters resulted in a single label being produced for all pairs, we performed another round of hyperparameter tuning, considering learning rate=[2e-8, 2e-7, 2e-5, 2e-5, 2e-4, 2e-3, 2e-2] and weight decay=[0, 1e-5, 1e-3, 1e-2, 1e-1], where the best-performing hyperparameters were learning rate=2e-4 and weight decay=1e-5.
 P1.
-For P1, we re-implemented the approach following the settings provided by the authors.
+For P1, we reimplemented the approach using the authors’ settings.
 The prompt was executed on GPT-4o with a temperature of 0 and a random seed of 133742243.
@@
 P2.
-For P2, we re-implemented the approach following the settings provided by the authors.
+For P2, we reimplemented the approach using the authors’ settings.
 While the original paper used GPT-3.5-turbo, our comparative experiments revealed that GPT-4o yielded better outcomes.
-Therefore, the results for P2 presented in this paper are based the GPT-4o model.
-Moreover, the choice of random seed and temperature followed the best-performing settings reported in their paper.
-P3. For P3_1 and P3_2, we re-implemented the approach using the settings provided by the authors over GPT-4o.
+Therefore, the results for P2 presented in this paper are based on the GPT-4o model.
+Moreover, the choice of random seed and temperature was based on the best-performing settings reported in their paper.
+P3. For P3_1 and P3_2, we reimplemented the approach using the authors’ settings on GPT-4o.
 4.5 Pre-trained Model Selection (RQ1)
@@
 We shortlist the ST models for investigation in our work based on the NLP leaderboard, which reports the 38 most accurate pre-trained models [^8] https://www.sbert.net/docs/pretrained_models.html.
-These models have been extensively evaluated for their ability to generate sentence embeddings (i.e., capturing the semantics of the whole text) and their performance in semantic search (i.e., finding relevant answers to a given query).
+These models have been extensively evaluated for their ability to generate sentence embeddings (i.e., capturing the semantics of the entire text) and for their performance in semantic search (i.e., retrieving relevant answers to a given query).
 Both tasks closely align with our objectives.
@@
 FPR is the proportion of negatives incorrectly identified as positives (i.e., the percentage of trace links wrongly identified as not trace links).
-The AUC of the ROC curve (computed as micro-average over all the provisions to avoid the dominance of some provisions) provides a single aggregate performance measure across all possible thresholds and, hence, is a suitable evaluation metric to compare the ST models.
-We posit that the model with the highest AUC value demonstrates the best overall performance in identifying trace links in a zero-shot setting, as a higher AUC value indicates a better balance between correctly identifying true trace links (high TPR) and minimizing the identification of false links (low FPR).
+The AUC of the ROC curve (computed as a micro-average across all provisions to avoid the dominance of any single provision) provides a single aggregate performance measure across all possible thresholds and, hence, is a suitable evaluation metric for comparing ST models.
+We posit that the model with the highest AUC performs best in identifying trace links in a zero-shot setting, as a higher AUC indicates a better balance between correctly identifying true trace links (high TPR) and minimizing false positives (low FPR).
 Results.
@@
 The best-performing model on HIPAA is ST29 ($K^{\dagger}=1$), with an AUC value of 0.859.
-The next best performing model is ST21 with an AUC value of 0.850.
+The next best-performing model is ST21, with an AUC of 0.850.
 The difference between these two AUC values is only marginal.
@@
 ST29 has been, however, trained on more (multi-lingual) data.
-Additionally, we observe a discrepancy in the performance of the models on the HIPAA dataset compared to that reported by the NLP community.
-The best NLP model, ST1, does not perform well on HIPAA, ranked 16.
-This observation indicates that well-performing models in NLP are not necessarily as effective for RE-specific problems.
-We acknowledge that zero-shot performance does not always predict fine-tuning effectiveness, particularly across different model architectures.
+Additionally, we observe a discrepancy between the models’ performance on the HIPAA dataset and that reported by the NLP community.
+The best NLP model, ST1, does not perform well on HIPAA, ranking 16th.
+This observation indicates that high-performing NLP models are not necessarily effective for RE-specific problems.
+We acknowledge that zero-shot performance does not always predict the effectiveness of fine-tuning, particularly across different model architectures.
 To address this concern, we fine-tuned the top five models based on their zero-shot AUC scores to evaluate their performance more comprehensively.
-However, our experiments show that ST29 still consistently outperforms the other models after fine-tuning.
+However, our experiments show that ST29 continues to outperform the other models after fine-tuning.
 While resource constraints prevented us from fine-tuning all possible models, this targeted evaluation provides evidence that ST29 is the most effective choice among the top-performing candidates.
@@
 We answer RQ2 on the benchmark dataset, HIPAA.
-Since HIPAA contains 10 requirements documents, we apply the leave-one-out (LOO) evaluation method, where $\mathcal{K}$ashif and baselines are tested on a left-out document each time and trained (or fine-tuned) on the remaining documents to emulate realistic settings.
-However, to ensure a reasonable balance between the training and test sets, we exclude one document (CCHIT, labeled H10 in Table 1) from the LOO process, as it contains 1,064 requirements, thereby including more than half the dataset.
+Since HIPAA comprises 10 requirements documents, we apply the leave-one-out (LOO) evaluation method, in which $\mathcal{K}$ashif and baselines are tested on a left-out document each time and trained (or fine-tuned) on the remaining documents to emulate realistic settings.
+However, to ensure a reasonable balance between the training and test sets, we exclude one document (CCHIT, labeled H10 in Table 1) from the LOO process because it contains 1,064 requirements, thereby comprising more than half of the dataset.
 Evaluation Metrics.
@@
 We report the mean and standard deviation across the nine documents.
-To ease comparisons of the models’ rankings, we also include Mean Average Precision (MAP) over requirements that have at least one trace link to a provision, which measures how well a model ranks relevant items higher than irrelevant ones.
+To facilitate comparisons of the models’ rankings, we also report Mean Average Precision (MAP) over requirements with at least one trace link to a provision, which measures how well a model ranks relevant items higher than irrelevant ones.
 We also report the AUC value for the overall performance of the classifiers.
@@
 Note that the MAP score is computed only for requirements containing at least one true positive, since requirements without any trace links provide no informative contribution to the MAP.
-In contrast, AUC is applied to all requirements as it allows for a more complete assessment of each classifier’s capability to distinguish relevant from non-relevant pairs, regardless of the presence or absence of TPs for a particular requirement.
+In contrast, AUC is applied to all requirements because it enables a more comprehensive assessment of each classifier’s ability to distinguish relevant from non-relevant pairs, regardless of the presence or absence of TPs for a particular requirement.
 Table 5: Accuracy of $\mathcal{K}$ashif and baselines on HIPAA (RQ2)
@@
 $\mathcal{K}$ashif $_{\text{tuned}}$ and the baseline models (LDA, LSI, B, and GloVe) require a decision threshold, which we tune using the same strategy explained in Section 3.
-In contrast, RoBERTa relies on the softmax function inherent to its architecture for final class-label assignment, and tuning the threshold over the output logits for label 1 resulted in a lower F2 score.
+In contrast, RoBERTa relies on the softmax function inherent to its architecture for final class-label assignment, and tuning the threshold on the output logits for label 1 yielded a lower F2 score.
 Similarly, the publicly available implementations of TraceBERT and LLaMA output both class probabilities and a final label, and using the label as the final output yielded a higher F2 score.
@@
 $\mathcal{K}$ashif $_{\text{tuned}}$ outperforms all variants of $\mathcal{K}$ashif in terms of precision, achieving an average of 68.1%.
-Based on our observations of the tuned thresholds, we found that most threshold values are greater than 0.5, which explains the higher precision compared to $\mathcal{K}$ashif $_{\text{constant}}$.
+Based on our observations of the tuned thresholds, we found that most threshold values exceed 0.5, which explains the higher precision relative to $\mathcal{K}$ashif $_{\text{constant}}$.
 For the H1 to H9 projects, the corresponding threshold values are 0.91, 0.5, 0.79, 0.46, 0.82, 0.81, 0.78, 0.73, and 0.58, respectively.
@@
 LDA achieves an F2 score of only 14.4% and a MAP of 37.8%, indicating its failure to assign higher ranks to true-positive pairs.
-Furthermore, its AUC of 0.51 suggests that LDA performs no better than random and offers little discriminatory power.
-In contrast, LSI attains an F2 score of 10.7%, which is 4 pp lower than LDA.
+Furthermore, its AUC of 0.51 indicates that LDA performs no better than chance and provides little discriminatory power.
+In contrast, LSI achieves an F2 score of 10.7%, which is 4 pp lower than that of LDA.
 However, it outperforms LDA in both MAP (53.5%) and AUC (0.71).
@@
 Therefore, the threshold-tuning process did not yield the best LSI performance.
-B, the ML-based technique tailored for the LRT task, exhibits relatively low retrieval performance compared to the other classifiers.
+B, an ML-based technique tailored to the LRT task, exhibits relatively low retrieval performance compared with the other classifiers.
 Nevertheless, it achieves the second-highest MAP score (78.8%), following $\mathcal{K}$ashif, indicating a strong ability to rank positive pairs higher than negative ones.
@@
 Its maximum attainable F2 score is approximately 22–23%, and the selected tuned thresholds achieve an F2 score of 22%.
-TraceBERT, however, performs poorly on the LRT task, with results comparable to LDA.
+TraceBERT, however, performs poorly on the LRT task, yielding results comparable to those of LDA.
 It achieves an F2 score of 8.7% and an AUC of 0.67, results that are only marginally better than random.
@@
 Its high AUC value indicates that RoBERTa effectively separates positive and negative pairs.
-However, its ranking quality shows greater variability than $\mathcal{K}$ashif and B, as reflected in its MAP, which is fourth-highest overall.
+However, its ranking quality shows greater variability than that of $\mathcal{K}$ashif and B, as reflected in its MAP, which is the fourth-highest overall.
 RoBERTa exhibits an almost constant F2 score across different threshold values, as shown in Figure 5(b).
@@
 Among the baselines, RoBERTa achieves the highest F2 score.
-RoBERTa also has the highest recall compared to all models and $\mathcal{K}$ashif variations, except for $\mathcal{K}$ashif Δ.
+RoBERTa also achieves the highest recall among all models and $\mathcal{K}$ashif variations, except for $\mathcal{K}$ashif Δ.
 However, compared to $\mathcal{K}$ashif constant, it has a lower recall, which shows weaker effectiveness in distinguishing relevant links by retrieving many FPs (650 FP links more than $\mathcal{K}$ashif constant).
@@
 LLaMa-w/o has the AUC value of 0.50, which shows a random behavior and lacks discrimination on the LRT task.
-However, LLaMa-tuned achieves an F2 score of 22.5%, an AUC of 0.80, and a MAP of 41.4%, indicating that while the model can distinguish between positive and negative pairs (as reflected in its relatively high AUC), its overall ranking and threshold-based performance remain limited.
+However, LLaMa-tuned achieves an F2 score of 22.5%, an AUC of 0.80, and a MAP of 41.4%, indicating that, although the model can distinguish positive from negative pairs (as reflected in its relatively high AUC), its overall ranking and threshold-based performance remain limited.
 The low MAP value suggests that LLaMa-tuned struggles to consistently assign high similarity scores to all true-positive pairs, capturing only a subset of links while missing many others.
@@
 Such a method and recall value can indeed be useful when building recommendation systems.
-However, they come at the cost of introducing more FPs (as evidenced by the low precision), requiring the human analyst to filter them out.
+However, they come at the cost of increasing FPs (as evidenced by low precision), requiring the human analyst to filter them out.
 Consequently, we select $\mathcal{K}$ashif $_{\text{constant}}$ as the best performing model for LRT.
@@
 We have conducted several experiments around this hypothesis.
-While we observed fewer FPs with a binary classifier, the overall improvement was not significant; hence, we do not report it in this paper.
+Although we observed fewer FPs with a binary classifier, the overall improvement was not statistically significant; therefore, we do not report it in this paper.
   • $\bullet$
@@
 Specifically, $\mathcal{K}$ashif achieves an F2 score of $\approx$63%.
-Compared to the best existing baseline from the literature, $\mathcal{K}$ashif has a gain of about 34 pp in F2 score.
+Compared with the best existing baseline in the literature, $\mathcal{K}$ashif achieves a gain of approximately 34 points in F2 score.
 Moreover, $\mathcal{K}$ashif achieves the highest MAP score of 0.81 and AUC value of 0.93 compared to the baseline classifiers.
@@
 In real-life scenarios, dealing with LRT involves navigating through many provisions, usually significantly more than 10, as in the simple HIPAA case.
-This inherent complexity is notable with the 26 provisions pertinent to software in the GDPR.
+This inherent complexity is notable with the 26 provisions pertaining to software in the GDPR.
 Using the test documents described in Section 4.2, we evaluate and compare two models, namely ST29—the best pre-trained ST model selected in RQ1 and $\mathcal{K}$ashif $_{\text{constant}}$—the best $\mathcal{K}$ashif variant fine-tuned on HIPAA identified in RQ2.
 Note that we opted not to fine-tune $\mathcal{K}$ashif again on the new documents for three reasons.
-First, the documents are small and thus inadequate for meaningful training (or fine-tuning).
-Second, we aim to challenge existing solutions with a more realistic scenario and apply them to new, unseen documents.
-Finally, $\mathcal{K}$ashif is a similarity-based solution which has been exposed to both the LRT task as well as the regulatory domain (terminology) in the first fine-tuning on HIPAA.
-Therefore, another fine-tuning is less likely to have any additional value.
+First, the documents are too small to support meaningful training (or fine-tuning).
+Second, we aim to evaluate existing solutions in a more realistic scenario and apply them to unseen documents.
+Finally, $\mathcal{K}$ashif is a similarity-based solution that has been exposed to both the LRT task and the regulatory domain (terminology) during the first fine-tuning on HIPAA.
+Therefore, further fine-tuning is unlikely to add value.
 Evaluation Metrics. To evaluate the models’ effectiveness, we report results at the requirements and trace-link levels.
@@
 Following this, we compute the success rate as the ratio of requirements for which the approach predicts correct trace links (including both partial and exact matches) to the total number of requirements.
-We also want to evaluate, on average, how well the model performs for a single requirement.
+We also aim to evaluate, on average, the model’s performance for a single requirement.
 Thus, we report two additional metrics: (iv) Macro Recall (MR), presented in Equation 1, to assess how well the predicted provisions are actually traced to the requirement; (v) Cost, presented in Equation 2, to capture the effort required for an analyst to check the predicted results and filter out the FPs.
@@
 A lower Cost corresponds to more efficient predictions, thereby reducing analyst effort.
-At the trace link level, we report the total number of actual trace links, true positives (TP) and false positives (FP), recall (R), precision (P), and F2 score.
+At the trace-link level, we report the total number of trace links, true positives (TPs), false positives (FPs), recall (R), precision (P), and F2 score.
 Table 7: Accuracy of ST29 and $\mathcal{K}$ashif on the test documents (RQ3).
@@
 Moreover, ST29 has the Cost of 0.06% and a MR of 1.3% per requirement.
-This means that for a requirement, on average, the model retrieves 0.06% of the provisions while ensuring 1.3% of the actual trace links are presented in those retrieved links.
+This means that, on average, for a given requirement, the model retrieves 0.06% of the provisions while ensuring that 1.3% of the actual trace links are included among those retrieved.
 $\mathcal{K}$ashif achieves a Cost of 4.7% while with the MR value of 12.4%.
-Fine-tuning the models improves performance on the test documents; however, the patterns, which are more complex, in the unseen test documents, are still not fully captured by $\mathcal{K}$ashif.
+Fine-tuning the models improves performance on the test documents; however, the more complex patterns in the unseen test documents are still not fully captured by $\mathcal{K}$ashif.
 To summarize, $\mathcal{K}$ashif consistently outperforms ST29 across all documents, with a notable difference in the number of partial matches.
-Results thus suggest that fine-tuning pre-trained models on a dedicated dataset is indeed necessary for the model to learn about the LRT task.
+Results thus suggest that fine-tuning pre-trained models on a dedicated dataset is necessary for the models to learn the LRT task.
 However, while better than the pre-trained model, $\mathcal{K}$ashif shows the following limitations: 1) it does not provide a rationale behind selecting a trace link, except for the fact that semantic similarity exceeds a pre-defined threshold.
@@
 The answer to RQ3 is that $\mathcal{K}$ashif outperforms ST29, demonstrating that fine-tuning helps the model learn about the LRT task.
-However, the performance of $\mathcal{K}$ashif still shows significant room for improvement in unseen domains.
+However, kashif’s performance still shows significant room for improvement in unseen domains.
 4.8 Effectiveness of Large Language Model (RQ4)
 The baselines’ performance on the LRT task is extremely poor, highlighting the need for improvement (RQ2).
-When we attempted a more refined approach using the ST models ($\mathcal{K}$ashif), it performed better than the baselines but did not achieve satisfactory results on an unseen dataset (RQ3).
-This indicates that the ST models can partially address some of the issues inherent in the baseline approaches.
+When we adopted a more refined approach using the ST models ($\mathcal{K}$ashif), it outperformed the baselines but did not yield satisfactory results on an unseen dataset (RQ3).
+This indicates that ST models can partially address issues inherent in baseline approaches.
 However, the ST models lack the robustness needed to generalize effectively across unseen data, as discussed in RQ3.
 Given their promising results across many tasks AroraHH24 ; abs_2310_18648 , RQ4 aims to assess whether prompting LLMs with an engineered prompt offers a meaningful alternative for LRT and to compare its effectiveness against prompts that were not engineered.
-We posit that LLMs, with their pre-training across different domains, might significantly improve trace link recovery.
+We posit that LLMs, with their pretraining across diverse domains, may significantly improve trace link recovery.
 Metholodology.
@@
 We begin by evaluating Rice_LRT under two temperature values (0 and 0.5) and five different random seed values.
-Then, we compare its performance to the baseline prompts using their reported best settings.
+We then compare its performance with the baseline prompts using their reported best settings.
 Table 8: Rice_LRT performance across different seed and temperature values.
@@
 In contrast, increasing randomness by setting the temperature equal to 0.5 caused a drop in F2 score for a constant seed value.
-The highest drop was observed for seed 16, where the F2 score dropped from 61.1% to 56.9%.
+The largest drop was observed for seed 16, from 61.1% to 56.9%.
 Moreover, we observe that the model’s behavior is not consistent across test documents.
-For example, when using seed value 13374224, increasing the temperature from 0 to 0.5 results in a decrease in F2 score for RD1, RD2, and RD4, while it leads to an improvement for RD3.
+For example, when using a seed value of 13374224, increasing the temperature from 0 to 0.5 decreases the F2 score for RD1, RD2, and RD4, whereas it increases it for RD3.
 Nonetheless, we observe that Rice_LRT shows little variation across different seeds when the temperature is set to 0, suggesting that random variation in results is primarily due to higher temperature.
-Thus, we select the performance using temperature 0 and arbitrarily select seed 16 to compare against the baselines.
+Thus, we select the performance at temperature 0 and arbitrarily set the seed to 16 to compare against the baselines.
+Moreover, we evaluated Rice_LRT using different number of few-shot examples (3, 3, and 4) to assess the sensitivity of its performance to the degree of few-shot supervision.
+We randomly chose 2, 3, and 4 examples from the existing few-shot examples.
+This created three different few-shot sets for each document.
+We ran the model separately with each set and recorded the resulting F2 scores.
+Based on the results, the average F2 score across 2-, 3-, and 4-shot settings are 56.5% (-4.6 pp), 58.9% (-2.2 pp), and 59.5% (-1.6 pp), respectively.
+Nonetheless, Rice_LRT consistently outperforms the baseline approaches across all example counts.
 Table 9 shows the results of the Rice_LRT approach and the baseline prompts.
@@
 P1 exhibits a lower number of retrieved links (TP+FP) compared to Rice_LRT and P2.
-Based on our observations of the rationale generated by the model, this stems primarily from its inability to capture indirect relationships between requirements and their corresponding provisions.
+Based on our observations of the model’s rationale, this stems primarily from its inability to capture indirect relationships between requirements and their corresponding provisions.
 For instance, a requirement stating that a user’s profile must include at least one charging point may seem purely functional but implicitly involves handling personal data, such as user identity or location.
-Consequently, it should be linked to the GDPR provision on obtaining user consent for data processing.
+Consequently, it should be linked to the GDPR provision governing the obtaining of user consent for data processing.
 The prompt’s inability to capture this implicit connection results in lower retrieval performance and the omission of true trace links.
@@
 P2 exhibits better performance, achieving a higher retrieval compared to P1.
-We observed that the rationales generated by P2 provide more details in explaining how to link specific requirements to the relevant provisions.
+We observed that the rationales generated by P2 provide more detail on how to link specific requirements to the relevant provisions.
 While P2’s high retrieval rate increased the number of TPs, it concurrently increased the FPs.
@@
 P3_1 is similar to P2 in terms of a low retrieval rate, while P3_2 resulted in producing the highest FPs compared to all other prompt templates.
-Moreover, $\mathcal{K}$ashif has a higher recall compared to P3_1 and P3_2 has a lower F2 score compared to $\mathcal{K}$ashif.
+Moreover, $\mathcal{K}$ashif has a higher recall compared to P3_1, and P3_2 has a lower F2 score compared to $\mathcal{K}$ashif.
 On the other hand, the inclusion of context and detailed instructions in Rice_LRT yields more effective outputs from GPT-4o, leading to improved performance at both the trace link and requirement levels.
@@
 Despite this, the number of partial matches has increased significantly, thereby improving the overall success rate.
-While one would ideally like an approach with a high exact-match rate, we note that the results remain beneficial, as we discuss next.
+While an approach with a high exact-match rate would be ideal, the results remain beneficial, as discussed next.
 Fig.
@@
 As seen in the figure, across all four documents, there were very few requirements with four FPs.
-This indicates that most partially matched requirements had a manageable number of FPs, typically ranging from one to three.
+This indicates that most partially matched requirements had a manageable number of FPs, typically between 1 and 3.
 This result is significant because it suggests that the model’s outputs are not overwhelming for analysts to process.
-Fewer FPs per requirement allow analysts to efficiently review and validate the suggested trace links, reducing their cognitive load.
+Fewer FPs per requirement allow analysts to review and validate the proposed trace links more efficiently, reducing their cognitive load.
 Instead of starting from scratch or sifting through a vast set of 26 possible provisions per requirement, analysts can focus on validating and refining a much smaller, pre-filtered set of trace links.
-This aligns with the principle of assisted decision-making skitka1999does , in which automated tools augment human judgment by narrowing options.
+This aligns with the principle of assisted decision-making skitka1999does , in which automated tools augment human judgment by narrowing the set of options.
 Our results further indicate that the GPT-4o model successfully demonstrated a stronger understanding of the LRT task when provided with an engineered prompt.
@@
 Its ability to navigate complex relationships and extract logical links demonstrates its robustness in understanding the nuances of regulatory requirements.
-However, the cases it misses highlight areas where the connections may require deeper domain-specific knowledge or additional context to resolve ambiguities.
+However, the cases it misses highlight areas in which the connections may require deeper, domain-specific knowledge or additional context to resolve ambiguities.
 On investigating the FPs for each requirement, we observed that several predicted trace links may be relevant depending on the application context, even though they do not exactly match the ground truth.
-These false positives include provisions that are not in the ground truth but are relevant to the input requirements, or provisions in the ground truth (but not all of them), which are why they are considered partial matches.
-This underscores the potential of Rice_LRT to identify trace links that correspond to potential associations between requirements and provisions that may not have been contemplated when building the ground truth.
+These false positives include provisions that are not in the ground truth but are relevant to the input requirements, as well as provisions in the ground truth (though not all of them), which is why they are considered partial matches.
+This underscores the potential of Rice_LRT to identify trace links that indicate associations between requirements and provisions that may not have been considered during ground-truth construction.
 Such cases could still be informative to the analysts.
@@
 The comparison between Rice_LRT and P1, , P3_1, and P3_2 yields values comparable to those reported for $\mathcal{K}$ashif, while the results for P2 are presented as follows: 0 and 5e-1, 1e-4 and 6e-1, 1e-4 and 5e-2, and 1e-4 and 3e-2 for RD1, RD2, RD3, and RD4, respectively.
-As we can see, across the four test documents, the resulting p-value is very small (below 5e-2).
-It falls well below the usual significance levels, indicating that the differences in precision and recall among $\mathcal{K}$ashif, P1, P2, P3_1, , P3_2 and Rice_LRT are statistically significant in most cases, except for P2 and the recall scores for RD1 to RD2.
+Across the four test documents, the resulting p-value is very small (below 5e-2).
+It falls well below the usual significance levels, indicating that the differences in precision and recall among $\mathcal{K}$ashif, P1, P2, P3_1, P3_2, and Rice_LRT are statistically significant in most cases, except for P2 and the recall scores for RD1 to RD2.
 The answer to RQ4 is that our Rice-based approach, which utilizes an engineered prompting on GPT-4o, outperforms the baselines on the LRT task across the four test documents.
-Further, the engineered context and instructions used in the few-shot setting of Rice_LRT achieved the highest F2 score by maximizing TPs while reducing FPs compared to the baselines.
+Furthermore, the engineered context and instructions used in the few-shot setting of Rice_LRT achieved the highest F2 score by maximizing TPs while minimizing FPs relative to the baselines.
 It also significantly outperforms $\mathcal{K}$ashif, although $\mathcal{K}$ashif remains competitive against P1 and P3_2 prompts.
@@
 We compared them against our ST-based classifier, $\mathcal{K}$ashif, on the LRT task using the publicly available HIPAA dataset.
-The results show that $\mathcal{K}$ashif achieves the highest F2 score among all baselines, including B, TraceBERT, and LLaMA, which have been previously proposed for inner-requirements traceability.
+The results show that $\mathcal{K}$ashif achieves the highest F2 score among all baselines, including B, TraceBERT, and LLaMA, which have been previously proposed for inter-requirements traceability.
 Moreover, $\mathcal{K}$ashif achieved a strong overall ranking performance, obtaining an MAP of 81% and the highest discrimination across thresholds, with an AUC of 0.93.
@@
 Rice_LRT achieves an MR of 84% with a Cost of 9.5%, meaning that, on average, only 9.5% of the total provisions are retrieved, while ensuring that 84% of the true trace links are included in the retrieved list.
-This represents a 36 pp improvement in MR over the best baseline prompt, while reducing Cost by 13 pp.
+This represents a 36-pp improvement in MR over the best baseline prompt, while reducing Cost by 13 pp.
 Moreover, we show that simple Yes/No–style prompts achieve low MR and Cost retrieval on the test documents, indicating that they do not fully leverage GPT-4o’s capabilities.
@@
 Bias is a well-known internal validity concern.
-To mitigate bias, in RQ3 and RQ4, the dataset of over four documents was curated by two annotators with more than a decade of experience in RE.
+To mitigate bias, for RQ3 and RQ4, the dataset comprising more than four documents was curated by two annotators with more than a decade of experience in RE.
 Before the traceability sessions, there was no exposure to the technical details of our approach.
@@
 The initial few-shot examples used for GPT-4o’s prompt engineering could introduce confirmation bias, potentially influencing the model’s predictions.
-To mitigate this, we designed the few-shot examples to reflect realistic usage scenarios where LLM is a recommendation tool guided by a human expert’s rationale for the first few requirements.
+To mitigate this, we designed the few-shot examples to reflect realistic usage scenarios in which an LLM serves as a recommendation tool guided by a human expert’s rationale for the first few requirements.
 This approach aligns with practical applications while minimizing the risk of confirmation bias.
 Additionally, the limited number of examples in the few-shot prompt was deliberately chosen to avoid overfitting.
-By doing so, we allowed the LLM sufficient flexibility to independently apply reasoning across the remaining requirements, maintaining a balance between guidance and adaptability.
+By doing so, we provided the LLM sufficient flexibility to independently apply reasoning to the remaining requirements, while maintaining a balance between guidance and adaptability.
 This approach ensures the LLM’s outputs remain broadly applicable while minimizing potential validity threats, as seen by the relatively high success rate in RQ4.
@@
 HIPAA is a pre-existing dataset frequently used in the RE literature.
-The test documents used in RQ3 and RQ4 (four new documents against GDPR), which we created as part of our work, covers two types of textual requirements, including user stories and shall-type requirements.
+The test documents used in RQ3 and RQ4 (four new documents against GDPR), which we created as part of our work, cover two types of textual requirements, including user stories and shall-type requirements.
 Such diversity helped increase the generalizability of our results.
 Experiments on more diverse requirements documents and other regulations are nonetheless required to improve the external validity of our study.
+Moreover, our evaluation is based on GPT-4o since this model is widely used in requirements engineering.
+Results obtained with other LLMs may differ; however, we expect them to be similar to those of LLMs with the same capabilities (focused on textual data and reasoning), and we leave a multi-model evaluation to future work.
 6 Related Work
 Requirements traceability (RT) has been extensively studied in RE mucha2024systematic ; li2023applications ; tufail2017systematic ; wang2018requirements ; ramesh1998factors .
-Existing work applies different technologies, ranging from traditional methods such as Information Retrieval (IR) and statistical models to more advanced approaches like Machine Learning (ML) and Deep Learning (DL).
+Existing work employs a range of technologies, from traditional methods such as Information Retrieval (IR) and statistical models to more advanced approaches such as Machine Learning (ML) and Deep Learning (DL).
 Early works borrowed IR techniques such as Vector Space Models (VSM), Latent Dirichlet Allocation (LDA), to find trace links between software artifacts via text relevancy gao2022using ; guo2017tackling ; kuang2017analyzing ; nishikawa2015recovering ; panichella2013and ; Mahmoud2013 ; Wang20193 ; sun2017frlink ; capobianco2013improving ; chhabra2017filtering ; bavota2014enhancing ; shao2013improved ; Wang20191 ; Li2020 ; moran2020improving .
 More advanced techniques have been introduced using ML Rasiman2022 ; Bella2019 ; mills2019tracing ; Bella2018 ; Mills2018 ; Rath2018 ; Zhao2018 ; Mills2017 ; Mills20172 ; Falessi2017 ; Hayes2015 ; le2015rclinker ; li2015recovering and DL kenton2019bert ; Zhang2021 ; Wang20192 ; Guo:17 ; Chen2019 ; Alhoshan2019 ; Alhoshan2018 ; Li2018 ; zhao2020extended ; Sultanov2013 , employing various algorithms — from classifiers like SVM, random forest, and decision trees to more sophisticated language models like BERT devlin-etal-2019-bert to find trace links.
-In recent years, with the emergence of LLMs, researchers have leveraged pre-trained knowledge through prompt engineering techniques to identify trace links between software artifacts hassine2024llm ; rodriguez2023prompts ; vogelsang2025impact .
+In recent years, with the emergence of LLMs, researchers have leveraged pre-trained knowledge via prompt engineering to identify trace links among software artifacts hassine2024llm ; rodriguez2023prompts ; vogelsang2025impact .
 Hassine hassine2024llm proposed an LLM-based technique that uses zero-shot learning on GPT3.5 to find trace links between requirements and goals in Goal-oriented Language (GRL) models.
 Moreover, Rodriguez et al.
-rodriguez2023prompts proposed an approach that integrates zero-shot prompting with reasoning to enhance results in the Traceability Link Recovery (TLR) problem on diverse software artifacts.
+rodriguez2023prompts proposed an approach that integrates zero-shot prompting with reasoning to improve performance on the Traceability Link Recovery (TLR) problem across diverse software artifacts.
 They have shown that a prompt that performs well with one model or dataset may not yield optimal results with another, underscoring the need to tailor prompts to the specific context.
@@
 hey2025requirements used Retrieval-Augmented Generation (RAG) with LLMs for inter-requirements traceability, where the model first retrieves relevant candidate requirements.
-Then, it identifies trace links among the retrieved candidates using the KISS prompt, which consists of a yes/no question prompting the model to determine whether a trace link exists between a given pair of artifacts.
+Then, it identifies trace links among the retrieved candidates using the KISS prompt, which asks the model whether a trace link exists between a given pair of artifacts.
 Fuchß fuchss2025beyond et al.
@@
 Additionally, Ronanki et al.
-ronanki2024requirements developed and experimented with five distinct interactive prompting patterns, each tuned for a specific dataset.
+ronanki2024requirements developed and evaluated five distinct interactive prompting patterns, each tuned to a specific dataset.
 In addition to the algorithms being used, the types of artifacts with which these algorithms are intended to work also play a significant role.
@@
 Cleland-Huang et al.
-cleland:2010 propose a probabilistic approach that identifies trace links between requirements and the HIPAA regulation by computing probability values based on detecting indicator terms for regulations that correspond to requirements.
+cleland:2010 propose a probabilistic approach that identifies trace links between requirements and the HIPAA regulation by computing probabilities based on the detection of indicator terms for regulations that correspond to requirements.
 The authors further propose extending the indicator terms with more domain-specific terms retrieved from the web.
-In a follow-up work, Gibiec et al.
-Gibiec:2010 further investigate mining the web.
+In follow-up work, Gibiec et al.
+Gibiec:2010 further investigate web mining.
 Guo et al.
 Guo:17 extend the previous two papers to improve the terminology gap problem, i.e., the mismatch between terms in requirements and regulations.
-The authors investigate different methods based on classification, ontologies, and web-mining and evaluate their approaches on HIPAA.
+The authors investigate methods based on classification, ontologies, and web mining and evaluate them on HIPAA.
 While previous research has made significant strides in requirements traceability using traditional IR methods and ML/DL techniques, these approaches exhibit notable limitations in addressing the complexities of the LRT task.
 Most notably, existing methods struggle with the terminology gap between regulations and technical requirements, do not generalize well across regulations, and lack adaptability to multi-domain applications.
-Additionally, studies that have proposed prompt-based approaches have not fully leveraged the capabilities of prompt engineering, providing only detailed instructions and limited context.
-Their approaches primarily rely on pre- and post-processing strategies hey2025requirements ; fuchss2025beyond , which in some cases query several large language models.
+Additionally, studies proposing prompt-based approaches have not fully leveraged the capabilities of prompt engineering, providing only detailed instructions and limited context.
+Their approaches primarily rely on pre- and post-processing strategies hey2025requirements ; fuchss2025beyond , which, in some cases, query multiple large language models.
 In comparison to the above works, we empirically evaluate two automated LRT approaches: (1) a classifier-based solution leveraging sentence transformers and (2) a generative LLM-based solution guided by structured prompt engineering.
@@
 To the best of our knowledge, we are also among the first to identify the strengths and limitations of LLMs in this context.
-Further and larger studies with human experts are required to establish the benefits of LLMs for LRT.
+Further, larger studies involving human experts are required to establish the benefits of LLMs for LRT.
 7 Conclusion
-This study presents a comparative evaluation of two approaches for Legal Requirements Traceability (LRT): a classifier-based method, $\mathcal{K}$ashif, leveraging sentence transformers, and a generative LLM-based method, Rice_LRT, designed using a structured prompt engineering framework.
+This study presents a comparative evaluation of two approaches to Legal Requirements Traceability (LRT): a classifier-based method, $\mathcal{K}$ashif, leveraging Sentence Transformers, and a generative LLM-based method, Rice_LRT, designed using a structured prompt engineering framework.
 Moreover, we compare $\mathcal{K}$ashif against seven distinct baselines—ranging from traditional IR techniques to recent deep learning and transformer-based models (LSI, LDA, GloVe, TraceBERT, RoBERTa, and LLaMa2)—to comprehensively demonstrate its capabilities.
@@
 Conversely, the Rice_LRT approach, built on generative LLMs, outperformed $\mathcal{K}$ashif on GDPR data, achieving an F2 score of 61% and reducing manual effort for traceability by enabling analysts to vet only a fraction of trace links.
-Moreover, Rice_LRT performed better than a simpler prompt, increasing TPs while reducing FPs.
+Moreover, Rice_LRT outperformed a simpler prompt, increasing TPs while reducing FPs.
 These findings suggest that generative LLMs and carefully designed prompts offer a promising approach to automating LRT tasks in complex legal domains.
@@
 We further plan to enhance LLM performance by incorporating domain-specific knowledge to better address terminology and contextual gaps between regulatory texts and technical requirements, particularly in the GDPR context.
-Acknowledgements
-The experiments conducted in this work were enabled by support provided by the Digital Research Alliance of Canada.
-Declarations
-Authors Contributions Romina Etezadi: Conceptualization, Investigation, Formal analysis, Software, Writing - original draft.
-Sallam Abualhaija: Conceptualization, Investigation, Data Curation, Formal analysis, Writing - original draft & review & editing.
-Chetan Arora: Conceptualization, Investigation, Data Curation, Formal analysis, Writing - review & editing.
-Lionel Briand: Conceptualization, Investigation, Formal analysis, Supervision, Writing - review & editing, Funding acquisition.
-Funding This work is supported by the H2020 COSMOS European project, grant agreement No.
-957254, NSERC of Canada under the Discovery and CRC programs, the Research Ireland grant 13/RC/2094-2, and the Luxembourg National Research Fund under grant number C23/IS/17958091/PLAITO.
+Acknowledgement.
+This work is supported by the H2020 COSMOS European project, grant agreement No.
+957254, NSERC of Canada under the Discovery and CRC programs, the Research Ireland grant 13/RC/2094-2, the Luxembourg National Research Fund under grant numbers C23/IS/17958091/PLAITO and NCER22/IS/16570468/NCER-FT.
 It is part of a collaborative research program between the University of Ottawa’s Nanda laboratory and the SnT centre at the University of Luxembourg.
-Conflict of Interest The authors declared that they have no conflict of interest.
-Ethical Approval Not applicable.
-Informed Consent Not applicable.
-Clinical Trial Number Not applicable.
+Declaration
+Conflict of Interest. The authors declared that they have no conflict of interest.